### v3

# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

Abstract
In 1976, Lai constructed a nontrivial confidence sequence for the mean $\mu$ of a Gaussian distribution with unknown variance $\sigma^{2}$. Curiously, he employed both an improper (right Haar) mixture over $\sigma$ and an improper (flat) mixture over $\mu$. Here, we elaborate carefully on the details of his construction, which use generalized nonintegrable martingales and an extended Ville’s inequality. While this does yield a sequential t-test, it does not yield an “e-process” (due to the nonintegrability of his martingale). In this paper, we develop two new e-processes and confidence sequences for the same setting: one is a test martingale in a reduced filtration, while the other is an e-process in the canonical data filtration. These are respectively obtained by swapping Lai’s flat mixture for a Gaussian mixture, and swapping the right Haar mixture over $\sigma$ with the maximum likelihood estimate under the null, as done in universal inference. We also analyze the width of resulting confidence sequences, which have a curious polynomial dependence on the error probability $\alpha$ that we prove to be not only unavoidable, but (for universal inference) even better than the classical fixed-sample t-test. Numerical experiments are provided along the way to compare and contrast the various approaches, including some recent suboptimal ones.
1 Introduction
The classical location tests for Gaussian data fall into either of the two categories: the Z-test, where the population variance is assumed to be known a priori, and a sample average is rescaled by the sample deviation to create a standardized test statistic; or the t-test [Student, 1908], where the population variance is unknown, and a plug-in estimator of the variance is used in lieu of the variance, accounting for the heavier (“regularly varying”) tail of the t-test statistic than the Gaussian Z-test statistic.
This paper concerns sequential analogs of t-tests. Sequential Z-tests have been widely established since Wald [1945]. One-sided or “power-one” variants have even been generalized nonparametrically to subGaussian data [Robbins, 1970, Howard et al., 2021], and recently to any data in the square-integrable class [Wang and Ramdas, 2023b].
Sequential extensions of the t-test were studied by Rushton [1950], Hoel [1954], Ghosh [1960], Sacks [1965], who were interested in calculating the approximate thresholds to control the type 1 error at a particular predefined stopping time. Inspired by the sequential Z-test work of Robbins, Lai [1976] constructed a sequential t-test that controls a different type of error that will be of particular interest to us: nonasymptotic, conservative (i.e. non-approximate) time-uniform control of the type 1 error at all arbitrary stopping times, possibly not specified in advance. We will define this formally later, but we mention at the outset that the filtration will matter very much: some methods will have guarantees at any stopping times in the original data filtration, while others only have guarantees in a reduced filtration, and studying this phenomenon in the context of t-tests will be an important part of this paper.
Lai and more recent authors [Grünwald et al., 2023, Pérez-Ortiz et al., 2022] situate the Gaussian t-test problem into the broader framework of group-invariant tests; indeed, the t-test null is invariant under rescaling of observations by a constant. The tests are mostly constructed using variants of the sequential likelihood ratio test (LRT) by Wald [1945], utilizing the fact that the likelihood ratios are nonnegative martingales (see Section 2.5). Building upon these group invariance tools, Lindon et al. [2022] construct a sequential F-test for the regression set-up, which coincides with one of the t-tests that we shall present when no covariate exists. A slightly different approach is taken by Jourdan et al. [2023] who study the same problem in the bandit context, sequentially bounding the empirical mean and variance separately using existing Z-test-type bounds and obtaining via a union bound a sequence of confidence “boxes” for mean and variance simultaneously. We shall return to and compare our bounds with some of these results later.
We shall focus in this paper primarily on the construction of e-processes, which are sequences of test statistics whose realized values reflect the evidence against the null in favor of the alternative. These e-processes can be thresholded and/or inverted to yield nonasymptotic, conservative, “anytime-valid” hypothesis tests and confidence intervals (called “confidence sequences”), but we emphasize that they are worthy of being used in their own right as well. Indeed, these e-processes directly measure evidence against the null: values below one are expected under the null, and large(r) values indicate (more) evidence against the null.
Nonnegative supermartingales are the prototypical examples of e-processes [Ramdas et al., 2022]. We emphasize that we are testing a composite null hypothesis, and it is not apriori obvious how to construct such “simultaneous” or “composite” nonnegative supermartingales or e-processes, or to compare various options. We shall formalize these concepts in greater detail in Sections 2.2 and 2.3. The main contributions of this paper are listed as follows:
  1. 1.
    We apply the method of “Universal Inference”, a variant of the LRT by Wasserman et al. [2020], to construct a t-test e-process for the point null $\mu=\mu_{0}$, as well as for the one-sided composite null $\mu\leqslant\mu_{0}$, implying respectively a two-sided and a one-sided confidence sequence for the mean $\mu$, both of which are new.
  2. 2.
    We fill in the missing details of Lai’s [1976] sequential t-test, looking into the curious question of why his approach yields a confidence sequence without an e-process, using the theory of non-$L^{1}$ nonnegative supermartingales [Wang and Ramdas, 2023a].
  3. 3.
    We introduce a simple variant of Lai [1976], deriving instead closed-form e-processes for the null $\mu=\mu_{0}$. The resulting confidence sequence enjoys improved tightness, and unlike Lai’s original, does not require numerical inversion or thresholding.
  4. 4.
    We derive the information-theoretic lower bound on the width of confidence intervals, and upper bound on the growth of e-processes for the t-test, both of which are attained, or almost attained, by our results (Contributions 1 and 3 in this list). We also show that other conservative sequential t-tests available in the literature (Lai’s original method, the plug-in method by Jourdan et al. [2023], the median method by Howard and Ramdas [2022]), and very surprisingly, even the fixed-time classical t-test, fail to be optimal in these regards.
2 Preliminaries
2.1 Notations
A universal probability space $(\Omega,\mathcal{A},\mathbb{P})$ is used when we consider the randomness of data. We use the indexed letters $X_{1},X_{2},\dots$ to denote an infinite stream of observations, taking values in $\mathbb{R}$. We denote by $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ the canonical filtration generated by the data, i.e. $\mathcal{F}_{n}=\sigma(X_{1},\dots,X_{n})$. Our shorthand notations on the sample include the partial sum $S_{n}=\sum_{i=1}^{n}X_{i}$, the sample average $\widehat{\mu}_{n}=S_{n}/n$, the partial sum of squares $V_{n}=\sum_{i=1}^{n}X_{i}^{2}$, the average square $\overline{X^{2}_{n}}=V_{n}/n$, the sample variance without Bessel’s correction $v_{n}=V_{n}/n-S_{n}^{2}/n^{2}$, and the t-statistic $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu_{0}}{\sqrt{nV_{n}-S_{n}^{2}}}$ where $\mu_{0}$ is the actual mean of $X_{1},X_{2},\dots$.
Usually, we consider i.i.d. observations $X_{1},X_{2},\dots$; and we use italic $P$ etc. to denote distributions over $\mathbb{R}$, while calligraphic $\mathcal{P}$ etc. to denote classes of distributions. For sample-dependent events and random variables, we carefully adopt the notations $P(\,\cdot\,)$ as
$\mathbb{P}[\,\cdot\,]\quad\text{when }X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P,$ (1)
and $\mathrm{E}_{P}(\,\cdot\,)$ as
$\mathbb{E}[\,\cdot\,]\quad\text{when }X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P.$ (2)
The conditional expectation $\mathrm{E}_{P}(\,\cdot\,|\mathcal{F})$ given $\mathcal{F}\subseteq\mathcal{A}$ is defined similarly. Univariate normal distribution with mean $\mu$ and variance $\sigma^{2}$ will be referred to as $N_{\mu,\sigma^{2}}$ – so “$N_{0,1}(X_{1}>0)=0.5$” means “$\mathbb{P}[X_{1}>0]=0.5$ when $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}$ standard normal”; its density is denoted by $p_{\mu,\sigma^{2}}$. The class of all normal distributions on $\mathbb{R}$ (excluding the degenerate $\sigma=0$ case) is denoted by $\mathcal{N}$. Classes of normal distributions with specified mean (or range of means) but any positive variance are denoted by, e.g., $\mathcal{N}_{\mu=\mu_{0}}$, $\mathcal{N}_{\mu\leqslant\mu_{0}}$; those with specified standardized mean by, e.g., $\mathcal{N}_{\mu/\sigma=\theta_{0}}$.
[…]
2.2 Sequential Statistics
The classical sequential analysis traces back to Wald [1947], and is formulated as testing the null $H_{0}:P\in\mathcal{P}_{0}$ against the alternative $H_{1}:P\in\mathcal{P}_{1}$, making a decision (to reject or not) each time a new data point $X_{n}\sim P$ is seen until a rejection is made. The type 1 error $P(H_{0}\text{ is ever rejected})$ for $P\in\mathcal{P}_{0}$ is to be controlled under a prescribed $\alpha$. This leads to the definition of any-time valid p-value [Johari et al., 2022], as a stochastic process $\{p_{n}\}_{n\geqslant 1}$ such that $P(p_{\tau}\leqslant p)\leqslant p$ for any stopping time $\tau$, any $p\in(0,1)$, and any $P\in\mathcal{P}_{0}$. Reject at the hitting time $\inf\{n:p_{n}\leqslant\alpha\}$ clearly controls the type I error.
The classical duality between tests and interval estimates also manifests in the sequential setting. Let $\mathcal{P}$ be a family of distributions and let $\theta:\mathcal{P}\to\mathbb{R}$ be the parameter of interest. A $(1-\alpha)$-confidence sequence (CS) [Darling and Robbins, 1967] over $\mathcal{P}$ for $\theta$ is a sequence of confidence intervals $\{\operatorname{CI}_{n}\}_{n\geqslant 1}$ such that, for any $P\in\mathcal{P}$, $P(\forall n\geqslant 1,\;\theta(P)\in\operatorname{CI}_{n})\geqslant 1-\alpha$; or equivalently, $P(\theta(P)\in\operatorname{CI}_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$. If the null $\mathcal{P}_{0}\subseteq\mathcal{P}$ is defined as a preimage $\theta^{-1}(\theta_{0})$, a sequential test can be constructed via rejection $H_{0}$ whenever $\theta_{0}\notin\operatorname{CI}_{n}$, and vice versa.
In this paper, the asymptotic tightness of a confidence sequence is measured in terms of (1) its growth rate, the rate at which its width increases when $n$ and the sample are fixed while $\alpha\to 0$ (or while $\alpha\downarrow\alpha_{\star}(n)$ if $\alpha$ is not allowed to be below $\alpha_{\star}(n)$); and (2) its shrinkage rate, the rate at which its width decreases when $\alpha$ is fixed while $n\to\infty$). For example, it is known that CSs for the mean over all 1-subGaussian random variables, and over all random variables with variance $\leqslant 1$ can both attain the minimax rates of growth $\sqrt{\log(1/\alpha)}$ and shrinkage $\sqrt{\log\log(n)/n}$ [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, Wang and Ramdas, 2023b].
2.3 Test Supermartingales and Extensions
Sequential tests and confidence sequences are usually constructed via nonnegative supermartingales (NSMs) and more generally, e-processes [Howard et al., 2021, Ramdas et al., 2022]. Let us, following Ruf et al. [2023], define them and state their key properties in composite terms.
Consider a stochastic process $\{M_{n}\}_{n\geqslant 0}$ where each $M_{n}$ is defined as a $[0,\infty]$-valued function of $X_{1},\dots,X_{n}$. We say it is
  • a nonnegative supermartingale (NSM) for $P$ on a filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$ if $\mathrm{E}_{P}(M_{0})<\infty$, $M_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathrm{E}_{P}(M_{n+1}|\mathcal{G}_{n})\leqslant M_{n}$ for all $n\geqslant 0$;
  • a nonnegative martingale (NM) for $P$ if, further, the equality $\mathrm{E}_{P}(M_{n+1}|\mathcal{G}_{n})=M_{n}$ holds.
Recently, Wang and Ramdas [2023a] defined a wider class of processes by dropping the integrability assumption $\mathrm{E}_{P}(M_{0})<\infty$, as the conditional expectation $\mathrm{E}_{P}(M_{n+1}|\mathcal{G}_{n})$ can still be well-defined as long as $M_{n+1}$ is nonnegative. We say the process $\{M_{n}\}$ is
  • an extended nonnegative supermartingale (ENSM) for $P$ on $\{\mathcal{G}_{n}\}$ if $M_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathrm{E}_{P}(M_{n+1}|\mathcal{G}_{n})\leqslant M_{n}$ are satisfied, irrespective of the finiteness of $\mathrm{E}_{P}(M_{0})$.
Recalling that $\mathcal{P}$ denotes a class of distributions, if, for all $P\in\mathcal{P}$, the process $\{M_{n}\}$ is NSM (or NM, ENSM) for $P$, we say it is an NSM (or NM, ENSM) for $\mathcal{P}$.
Further, we say $\{M_{n}\}$ is
  • an e-process for $P$ on $\{\mathcal{G}_{n}\}$, if $M_{n}$ is $\mathcal{G}_{n}$-measurable and, for each $P\in\mathcal{P}$, there is an NSM $\{N_{n}^{P}\}$ with $N_{0}^{P}=1$ on $\{\mathcal{G}_{n}\}$ for $P$ that upper bounds $\{M_{n}\}$: $P(M_{n}\leqslant N_{n}^{P})=1$.
Note that the filtration $\{\mathcal{G}_{n}\}$ is usually the canonical one $\{\mathcal{F}_{n}\}$, but as we shall see later, other non-trivial choices will lead to interesting results.
A parametrized family of NSMs or ENSMs can lead to a new NSM or ENSM by mixture. If $\{M_{n}(\theta)\}$ is an NSM for $\mathcal{P}$ for any $\theta\in\Theta$, so is the mixture $\{\int M_{n}(\theta)\mu(\mathrm{d}\theta)\}$ for any finite measure $\mu$ over $\Theta$, under mild measurability assumptions; if each is an ENSM for $\mathcal{P}$, the mixture is still an ENSM for $\mathcal{P}$ as long as $\mu$ is $\sigma$-finite [Wang and Ramdas, 2023a, Section 5.2]. Similar results hold straightforwardly for the mixtures of e-processes as well.
An important property of these processes is Ville’s inequality [Ville, 1939] and the recently discovered extended Ville’s inequality [Wang and Ramdas, 2023a, Theorem 4.1]. These maximal inequality bounds the crossing probability of a NSM (and hence e-process) or ENSM over an unbounded time horizon.
Lemma 2.1 (Ville’s inequality).
If $\{M_{n}\}_{n\geqslant 0}$ is an NSM for $\mathcal{P}$, for all $P\in\mathcal{P}$, $\varepsilon\in(0,1)$,
$P(\exists n\geqslant 0,\ M_{n}\geqslant\varepsilon)\leqslant\varepsilon^{-1}\mathrm{E}_{P}(M_{0}).$ (3)
Consequently, if $\{M_{n}\}_{n\geqslant 0}$ is an e-processes for $\mathcal{P}$, for all $P\in\mathcal{P}$, $\varepsilon\in(0,1)$,
$P(\exists n\geqslant 0,\ M_{n}\geqslant\varepsilon)\leqslant\varepsilon^{-1}.$ (4)
Lemma 2.2 (Extended Ville’s inequality).
Let $\{M_{n}\}_{n\geqslant 0}$ be an ENSM for $\mathcal{P}$. Then for all $P\in\mathcal{P}$, $\varepsilon\in(0,1)$,
$P(\exists n\geqslant 0,\ M_{n}\geqslant\varepsilon)\leqslant\varepsilon^{-1}\mathrm{E}_{P}(\mathbbmss{1}_{\{M_{0}<\varepsilon\}}M_{0})+P(M_{0}\geqslant\varepsilon).$ (5)
Lemmas 2.1 and 2.2 suggest that, if $\mathcal{P}_{0}$ is a set of null distributions, NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ can be seen as measurements of evidence against the null $H_{0}:P\in\mathcal{P}_{0}$. To see that, if we have at hand an NSM or e-process $\{M_{n}\}$ for $\mathcal{P}_{0}$ such that $M_{0}=1$, we may reject the null $\mathcal{P}_{0}$ whenever the process $M_{n}$ exceeds $1/\alpha$. This, due to Lemma 2.1, controls the sequential type 1 error $P(H_{0}\text{ is ever rejected})$ within $\alpha$ for any $P\in\mathcal{P}_{0}$. Similarly, if we have an ENSM $\{M_{n}\}$ for $\mathcal{P}_{0}$, we may set the right hand side of (5) to $\alpha$ and solve the corresponding $\varepsilon$, rejecting $H_{0}:P\in\mathcal{P}_{0}$ whenver $M_{n}$ exceeds $\varepsilon$. Desirably, an NSM, e-process or ENSM for $\mathcal{P}_{\theta_{0}}$ used as a powerful test process shrinks under the null $\mathcal{P}_{0}$ and grows under the alternative $\mathcal{P}_{1}$. A confidence sequence for a parameter $\theta$ follows in the same way if we have an NSM or an ENSM $\{M^{\theta_{0}}_{n}\}$ for every $\mathcal{P}_{\theta_{0}}=\{P\in\mathcal{P}:\theta(P)=\theta_{0}\}$: by solving the inequality $M_{n}^{\theta_{0}}\leqslant 1/\alpha$ or $M_{n}^{\theta_{0}}\leqslant\varepsilon$ in $\theta_{0}$, that is. An illustrative comparison between NSMs and ENSMs in the Gaussian case with known variance for both constructing tests and confidence sequences is provided recently by Wang and Ramdas [2023a, Section 5].
[…]
2.4 Sequential t-Test and t-Confidence Sequences
Different objectives are pursued for the sequential Gaussian t-test problem. Early authors have studied the problem of testing the null $\mathcal{N}_{\mu=0}$ (or more generally $\mathcal{N}_{\mu/\sigma=\theta_{0}}$) against the alternative $\mathcal{N}_{\mu/\sigma=\theta_{1}}$ [Rushton, 1950, Ghosh, 1960, Sacks, 1965], and its scale-invariant nature translates into the group-invariant setting as the problem of testing orbital nulls and alternatives [Pérez-Ortiz et al., 2022].
In this paper, we shall roughly follow the same routine as Lai [1976] and be less concerned with the alternatives, chiefly focusing instead on the construction of (1) test processes for the null $\mathcal{N}_{\mu=0}$ (and study their behavior under both null and non-null underlying distributions); (2) Confidence sequences for the population mean $\mu$ over the set of all Gaussian distributions $\mathcal{N}$, which we shall often shorthand as “t-confidence sequences”, or “t-CSs” for brevity. One may note that (1) is sufficient to produce tests for arbitrary point location null $\mathcal{N}_{\mu=\mu_{0}}$ by shifting, and consequently (2) by Lemmas 2.1 and 2.2 and inversion. Occasionally we are interested in test processes for the one-sided null $\mathcal{N}_{\mu\leqslant 0}$ as these give rise to one-sided tests and CSs. In fact, we shall later construct test processes for $\mathcal{N}_{\mu/\sigma=\theta_{0}}$ which is the strongest statistical quantity in the setting, as it leads to both CSs and tests for the mean $\mu$ and the normalized mean $\mu/\sigma$.
2.5 Likelihood Ratio Martingales
It is well known that the likelihood ratio process is a nonnegative martingale.
Lemma 2.3.
Let $Q\ll P$ be probability distributions on $\mathbb{R}$. Then, the process
$L_{n}=\prod_{i=1}^{n}\frac{\mathrm{d}Q}{\mathrm{d}P}(X_{i})$ (6)
is a nonnegative martingale for $P$ on the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
We shall use two stronger forms of Lemma 2.3, stated below and proved in Section C.1. First, the distribution on the numerator, $Q$, can be varying and depend on previous observations.
Lemma 2.4.
Let $P$ be a probability distribution on $\mathbb{R}$. For each $n\geqslant 1$ let $Q_{n}:\Omega\times\mathcal{B}(\mathbb{R})\to\mathbb{R}$ be a Markov kernel such that: 1) $Q_{n}(\,\cdot\,,B)$ is $\mathcal{F}_{n-1}$-measurable for all $B\in\mathcal{B}(\mathbb{R})$, and 2) $Q_{n}(\omega,\,\cdot\,)\ll P$ for all $\omega\in\Omega$. Then, the process
$L_{n}=\prod_{i=1}^{n}\frac{\mathrm{d}Q_{i}}{\mathrm{d}P}(X_{i})$ (7)
is a nonnegative martingale for $P$ on the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
[…]
Lemma 2.5.
[…]
3 t-Test e-Processes via Universal Inference
We shall demonstrate how the method of universal inference [Wasserman et al., 2020] leads to e-processes and confidence sequences for the t-test. Throughout this section, for $n\geqslant 0$, we denote by $\widetilde{\mu}_{n}$ and $\widetilde{\sigma}_{n}$ any point estimators for $\mu$ and $\sigma$ adapted to the canonical filtration $\mathcal{F}_{n}$ (i.e. based on $X_{1},\dots,X_{n}$). For example, they can simply be the sample mean $\widehat{\mu}_{n}$ and sample standard deviation $\sqrt{v_{n}}$; we may also use the Bayesian approach by sequentially updating them as the posterior mean from, say, a normal-inverse-gamma prior on $(\mu,\sigma)$111which is the conjugate prior of normal distributions when both of $\mu$ and $\sigma$ are unknown. The method of sequential universal inference [Wasserman et al., 2020, Section 8] implies the following plug-in likelihood ratio martingale for a fixed Gaussian distribution:
Corollary 3.1 (Universal inference Z-test martingale).
For any $\mu$, $\sigma$ and their point estimators $\{\widetilde{\mu}_{n}\}$, $\{\widetilde{\sigma}_{n}\}$ adapted to the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ (such as the sample mean and standard deviation), the process
$\ell_{n}^{\mu,\sigma^{2}}=\frac{\prod_{i=1}^{n}p_{\widetilde{\mu}_{i-1},\widetilde{\sigma}^{2}_{i-1}}(X_{i})}{\prod_{i=1}^{n}p_{\mu,\sigma^{2}}(X_{i})}=\frac{\sigma^{n}}{\prod_{i=1}^{n}\widetilde{\sigma}_{i-1}}\exp\left\{\sum_{i=1}^{n}\left(\frac{(X_{i}-\mu)^{2}}{2\sigma^{2}}-\frac{(X_{i}-\widetilde{\mu}_{i-1})^{2}}{2\widetilde{\sigma}^{2}_{i-1}}\right)\right\}$ (9)
is a martingale for $N_{\mu,\sigma^{2}}$ on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
It is not hard to see that Corollary 3.1 follows from Lemma 2.4, so we omit its proof. It is so named as the process $\{\ell_{n}^{\mu,\sigma}\}$ itself can be used, when $\sigma^{2}$ is known, as a sequential Z-test. From martingales for $N_{\mu,\sigma^{2}}$ for each $\mu$ and $\sigma$ to test processes for the t-test nulls, $\mathcal{N}_{\mu=0}$ and $\mathcal{N}_{\mu\leqslant 0}$, we can simply take the infima of (9) over $\{\mu=0,\sigma>0\}$ and $\{\mu\leqslant 0,\sigma>0\}$, which leads to e-processes for these nulls. They both have a closed-form expression and are stated below as Theorem 3.2 and Theorem 3.4, both proved in Section C.2.
Theorem 3.2 (Universal inference t-test e-process).
For any point estimators $\{\widetilde{\mu}_{n}\}$ and $\{\widetilde{\sigma}_{n}\}$ adapted to the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$, the process
$R_{n}=\left(\overline{X^{2}_{n}}\right)^{n/2}\mathrm{e}^{n/2}\prod_{i=1}^{n}\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}$ (10)
is an e-process for $\mathcal{N}_{\mu=0}$ on $\{\mathcal{F}_{n}\}$. Consequently, define
$W_{n}=\frac{1}{\alpha^{2/n}\mathrm{e}}\exp\left\{\frac{\sum_{i=1}^{n}\log\widetilde{\sigma}_{i-1}^{2}+\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}}{n}\right\}.$ (11)
Then $\operatorname{CI}_{n}=\left[\widehat{\mu}_{n}\pm\sqrt{\widehat{\mu}_{n}^{2}-\overline{X^{2}_{n}}+W_{n}}\right]$ forms a $(1-\alpha)$-CS for $\mu$ over $\mathcal{N}$.
It is worth remarking that for the general location null $\mathcal{N}_{\mu=\mu_{0}}$, one must resist the temptation to replace all $X_{i}$ in (10) above by $X_{i}-\mu_{0}$, which, while indeed produces an e-process for $\mathcal{N}_{\mu=\mu_{0}}$, loses its power. The correct modification is to only shift $X_{i}$ in the $\overline{X^{2}_{n}}$ term, i.e. using the e-process for $\mathcal{N}_{\mu=\mu_{0}}$
$R_{n}^{\mu_{0}}=\left(\frac{\sum_{i=1}^{n}(X_{i}-\mu_{0})^{2}}{n}\right)^{n/2}\mathrm{e}^{n/2}\prod_{i=1}^{n}\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}.$ (12)
This can be understood as shifting $X_{i}$ and also $\widetilde{\mu}_{i-1}$ by $\mu_{0}$. This is how we arrive at the confidence sequence $\operatorname{CI}_{n}=\left[\widehat{\mu}_{n}\pm\sqrt{\widehat{\mu}_{n}^{2}-\overline{X^{2}_{n}}+W_{n}}\right]$ above.
As a test process, the e-process $\{R_{n}\}$ distinguishes the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following asymptotic result on a limit which we refer to as the “growth index” of the test process.
Proposition 3.3 (Growth index of the universal inference t-test e-process).
Under any $P=N_{\mu,\sigma^{2}}$, suppose there is a $\gamma>0$ such that $\{\widetilde{\mu}_{n}\}$ converges to $\mu$ in $L^{3}$ with rate $\mathrm{E}_{P}(|\widetilde{\mu}_{n}-\mu|^{3})\lesssim n^{-\gamma}$, $\{\widetilde{\sigma}_{n}^{-2}\}$ converges to $\sigma^{-2}$ both in $L^{2}$ with rate $\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-2}-\sigma^{-2})^{2}\lesssim n^{-\gamma}$ and almost surely, and has uniformly bounded 3rd moment $\limsup_{n}\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-6})<\infty$. Then,
$\lim_{n\to\infty}\frac{\log R_{n}}{n}=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})\quad\text{almost surely}.$ (13)
Consequently, $\{R_{n}\}$ diverges almost surely to $R_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu\neq 0}$.
The assumptions on the point estimators are mild, as they are satisfied by the empirical mean and variance due to the moment properties of the inverse-$\chi^{2}$ distribution, as well as their smoothed or Bayesian extensions (posterior means under reasonable priors). The exponential growth of $\{R_{n}\}$ under the alternative $\mathcal{N}_{\mu\neq 0}$ is in contrast to its restrained behavior under the null $\mathcal{N}_{\mu=0}$, characterized by Ville’s inequality for e-processes (Lemma 2.1).
The $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ growth index in Proposition 3.3, we remark, is universal among test processes for t-tests. If one uses the plug-in likelihood ratio $\{\ell_{n}^{0,\sigma^{2}}\}$ in Corollary 3.1 to conduct the Z-test for the null $N_{0,\sigma^{2}}$, then, under the actual distribution $N_{\mu,\sigma^{2}}$, it is not hard to see that a faster rate $\frac{\log\ell_{n}^{0,\sigma^{2}}}{n}\to\mu^{2}/2\sigma^{2}$ holds under similar assumptions of the point estimators (which we show in Appendix B). We shall see in the rest of the paper many more occurrences of the same $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ limit, and we shall discuss that this is indeed the optimal rate in Section 6.3.
Apart from the asymptotics of the test process, we can also analyze the confidence sequence it implies. Let us briefly study the asymptotics of the radius $\sqrt{\widehat{\mu}_{n}^{2}-\overline{X^{2}_{n}}+W_{n}}=\sqrt{W_{n}-\left(\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)}$ of the CS in Theorem 3.2, as $\alpha\to 0$, $\sigma\to\infty$, and $n\to\infty$. Note that both $W_{n}$ and $\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}$ converge to $\sigma^{2}$, as long as the estimators $\tilde{\sigma}_{i}$ and $\tilde{\mu}_{i}$ are consistent. From (11), the deviation of $T_{n}$ from $\sigma^{2}$ is approximately $\sigma^{2}(\alpha^{-2/n}\exp(1/\sqrt{n})-1)$; While by the $\chi^{2}$-tail bound and the Chernoff bound,
$\mathbb{P}\left[\left|\frac{n}{n-1}\left(\overline{X^{2}_{n}}-{\widehat{\mu}_{n}}^{2}\right)-\sigma^{2}\right|\geqslant\sigma^{2}t\right]\lesssim\mathrm{e}^{-nt^{2}},$ (14)
so the deviation of $\overline{X^{2}_{n}}-{\widehat{\mu}_{n}}^{2}$ from $\sigma^{2}$ is approximately $\sigma^{2}/n$. Hence the radius of the CS scales, as a function of $\alpha$, $\sigma$ and $n$, at the rate of $\sigma\alpha^{-1/n}n^{-1/2}$. We remark that the “growth rate” of the CS, i.e., its dependence on $\alpha\to 0$, is in the time-dependent form of $\alpha^{-1/n}$ which is worse than the typical known-variance rate $\sqrt{\log(1/\alpha)}$. This, as we shall explain in Section 6.1, is also unavoidable.
Let us now state the e-process for the one-sided null.
Theorem 3.4 (Universal inference one-sided t-test e-process).
For any point estimators $\{\widetilde{\mu}_{n}\}$ and $\{\widetilde{\sigma}_{n}\}$ adapted to the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$, the process
$R^{-}_{n}=\left(\overline{X^{2}_{n}}-(\widehat{\mu}_{n}\wedge 0)^{2}\right)^{n/2}\mathrm{e}^{n/2}\prod_{i=1}^{n}\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}$ (15)
is an e-process for $\mathcal{N}_{\mu\leqslant 0}$ on $\{\mathcal{F}_{n}\}$.
There is clear similarity between the expression of the point null e-process (10), and that of the one-sided (15). The only difference is in that the former’s $\overline{X^{2}_{n}}$ term is here off by a $-(\widehat{\mu}_{n}\wedge 0)^{2}$ term in (15). When the data are predominantly negative, this term inhibits the exponential growth of the e-process. Therefore, while we shall witness exponential growth of (10) when the actual mean is negative (significant in the scale of standard deviation), no such growth is likely to happen in (15). We can formalize this by an asymptotic result similar to Proposition 3.3.
Proposition 3.5 (Growth index of the universal inference one-sided t-test e-process).
Under any $N_{\mu,\sigma^{2}}$, suppose the point estimators $\{\widetilde{\mu}_{n}\}$ and $\{\widetilde{\sigma}_{n}^{-2}\}$ satisfy the same assumptions as Proposition 3.3. Then,
$\lim_{n\to\infty}\frac{\log R_{n}^{-}}{n}=\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})\quad\text{almost surely}.$ (16)
Consequently, $\{R_{n}^{-}\}$ diverges almost surely to $R_{\infty}^{-}=\infty$ exponentially fast under $\mathcal{N}_{\mu>0}$.
A slightly weaker similarity between the expressions of these e-processes and the plug-in martingale (9) can also be observed under careful comparison, which we do in the “Test Process” row of Table 2, where we compare all aspects of universal inference methods for Z-test (including a one-sided test involving taking infimum over $t\leqslant 0$ in (9)) and t-test.
4 Sequential t-Tests via Scale Invariance
4.1 Lai’s [1976] Confidence Sequence
Let us first quote a theorem due to Lai [1976] who presented it almost without proof.
Theorem 4.1 (Lai’s t-Confidence Sequence; Theorem 2 of Lai [1976]).
Choose a starting time $m\geqslant 2$ and a constant $a>0$. Recall that $v_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}$. Further, define
$b$ $:=\frac{1}{m}\left(1+\frac{a^{2}}{m-1}\right)^{m},$ (17)
$\xi_{n}$ $:=\sqrt{v_{n}[(bn)^{1/n}-1]}.$ (18)
Then, the intervals $\operatorname{CI}_{n}=[\widehat{\mu}_{n}\pm\xi_{n}]$ satisfy, for any $\mu$ and $\sigma>0$,
$N_{\mu,\sigma^{2}}\left(\exists n\geqslant m,\mu\notin\operatorname{CI}_{n}\right)\leqslant 2(1-F_{m-1}(a)+af_{m-1}(a)),$ (19)
where $F_{m-1},f_{m-1}$ denote the CDF, PDF of t-distribution with $m-1$ degrees of freedom.
If we want $\{\operatorname{CI}_{n}\}$ to be a $(1-\alpha)$-CS over $\mathcal{N}$ for $\mu$, we need to solve $a$ from the equation $2(1-F_{m-1}(a)+af_{m-1}(a))=\alpha$ beforehand. To see the relationship between $a$ and $\alpha$, note that when $a$ is large,
$f_{m-1}(a)\asymp a^{-m},$ (20)
$1-F_{m-1}(a)\asymp a^{-(m-1)}.$ (21)
Hence
$\alpha\asymp a^{-(m-1)}\implies a\asymp\alpha^{-1/(m-1)}.$ (22)
The radius hence grows as (for fixed $m,n$ and $\alpha\to 0$)
$\xi_{n}\approx\sigma\sqrt{\left(\frac{1}{m}\left(1+\frac{a^{2}}{m-1}\right)^{m}n\right)^{1/n}-1}\asymp\alpha^{-\frac{m}{n(m-1)}}.$ (23)
In terms of its shrinkage rate (fixed $m,\alpha$ but $n\to\infty$), we have
$\xi_{n}\approx\sigma\sqrt{\frac{\log bn}{n}+\tilde{\mathcal{O}}(n^{-2})}.$ (24)
The CS is only valid from some time $m\geqslant 2$, which we shall soon explain (and, in some sense, remedy). The proof of this theorem, as it turns out, hinges on the extended Ville’s inequality (Lemma 2.2) for nonintegrable nonnegative supermartingale. Equally interestingly, it mixes a parametrized family of martingales under a coarser filtration. We shall begin our reworking of Lai’s CS from this concept, and eventually state and prove Theorem 4.7, a more revealing version of Theorem 4.1; as well as Theorem 4.9, a variant of Theorem 4.1 that does not involve an improper mixture and nonintegrability.
4.2 Scale Invariant Filtration
We say that a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is scale invariant if it is measurable and for any $x_{1},\dots,x_{n}\in\mathbb{R}$ and $\lambda>0$, $f(x_{1},\dots,x_{n})=f(\lambda x_{1},\dots,\lambda x_{n})$. Let us define the following sub-filtration of the canonical filtration $\{\mathcal{F}_{n}\}$.
Definition 4.2 (Scale invariant filtration).
For $n\geqslant 1$, let
$\mathcal{F}_{n}^{*}=\sigma(f(X_{1},\dots,X_{n}):f\text{ \emph{is scale invariant}}).$ (25)
Then, the filtration $\{\mathcal{F}^{*}_{n}\}_{n\geqslant 1}$ is called the scale invariant filtration of data $X_{1},X_{2},\dots$.
Definition 4.2 states that $\mathcal{F}^{*}_{n}$ is the coarsest $\sigma$-algebra to which all scale invariant functions of $X_{1},\dots,X_{n}$ are measurable. For example, recall that we denote by $T_{n-1}$ the t-statistic of the data $\sqrt{n-1}\frac{S_{n}-n\mu_{0}}{\sqrt{nV_{n}-S_{n}^{2}}}$, which is $\mathcal{F}^{*}_{n}$-measurable when $\mu_{0}=0$ and is a quantity frequently used to construct scale-invariant statistics later. To see that $\{\mathcal{F}^{*}_{n}\}$ is indeed a filtration, let $f$ be any scale invariant function $\mathbb{R}^{n}\to\mathbb{R}$, and define $g:\mathbb{R}^{n+1}\to\mathbb{R}$ as $g(x_{1},\dots,x_{n},x_{n+1})=f(x_{1},\dots,x_{n})$ which is also scale invariant. So $f(X_{1},\dots,X_{n})=g(X_{1},\dots,X_{n},X_{n+1})$ is $\mathcal{F}^{*}_{n+1}$-measurable. The arbitrariness of $f$ implies that $\mathcal{F}^{*}_{n}\subseteq\mathcal{F}^{*}_{n+1}$.
The filtration $\{\mathcal{F}_{n}^{*}\}$ contains all the information up to time $n$ about the relative sizes of the observations. For example, $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2X_{3}\}\in\mathcal{F}_{4}$ while $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2\}\notin\mathcal{F}_{4}$. The reader may recall the exchangeable sub-$\sigma$-algebra in the theory of exchangeability and backwards martingales such as in Klenke [2013, Chapter 12] for an analogy. Actually, when each of the observations $X_{1},X_{2},\dots$ is non-zero, it has the following clean expression.
Proposition 4.3.
If $X_{1}\neq 0$,
$\mathcal{F}_{n}^{*}=\sigma\left(\frac{X_{1}}{|X_{1}|},\frac{X_{2}}{|X_{1}|},\dots,\frac{X_{n}}{|X_{1}|}\right).$ (26)
Proof.
Let $\mathcal{F}_{n}^{**}=\sigma\left({X_{1}}/{|X_{1}|},{X_{2}}/{|X_{1}|},\dots,{X_{n}}/{|X_{1}|}\right)$. Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is scale invariant. Then, $f(X_{1},\dots,X_{n})=f(X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|)$, which is clearly $\mathcal{F}_{n}^{**}$-measurable, implying $\mathcal{F}_{n}^{*}\subseteq\mathcal{F}_{n}^{**}$. The inclusion $\mathcal{F}_{n}^{**}\subseteq\mathcal{F}_{n}^{*}$ is trivial: $x_{1}/|x_{1}|$, $x_{2}/|x_{1}|$, …, $x_{n}/|x_{1}|$ are themselves scale invariant functions of $x_{1},\dots,x_{n}$. ∎
In our scenario of application, the t-test under Gaussian distribution, data are not non-zero but almost surely non-zero. We do not distinguish (25) and (26) as the definition of the scale invariant filtration because processes adapted to it are often derived by manipulations of $X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|$, which we call the scale invariant reduction of the observations, ignoring the $N_{\mu,\sigma^{2}}$-negligible event $\{X_{1}=0\}$. Indeed, Lai [1976] defines such filtration via (26) for the t-test case. However, we remark that one would need the more general (25) in Definition 4.2 for the case when $X_{1}\neq 0$ does not hold almost surely (e.g. when exploring a nonparametric extension of our set-up and methods).
As a counter-example, one may verify that the e-processes for t-test that we have previously introduced via universal inference, $\{R_{n}\}$ in (10) and $\{R_{n}^{-}\}$ in (15), are not adapted to $\{\mathcal{F}_{n}^{*}\}$. They would have been if the plug-in estimates $\widetilde{\mu}_{i-1}$ and $\widetilde{\sigma}_{i-1}$ were allowed to be homogeneous of degree 1. However, as $\widetilde{\mu}_{0}$ and $\widetilde{\sigma}_{0}$ must be constants, this cannot be the case. In fact, when $n=1$, these two e-processes are necessarily non-constant, and thus are not $\mathcal{F}_{1}^{*}$-measurable.
4.3 Scale Invariant Likelihood Ratios
Let us suppose, temporarily in this subsection for the sake of generality, that $P_{\mu,\sigma^{2}}$ is a distribution on $\mathbb{R}$ parametrized by $\mu\in\mathbb{R}$ and $\sigma>0$, and $\eta$ a reference measure on $\mathbb{R}$, such that $P_{\mu,\sigma^{2}}\ll\eta$ with density
$\frac{\mathrm{d}P_{\mu,\sigma^{2}}}{\mathrm{d}\eta}(x)=\sigma^{-1}g(\sigma^{-1}(x-\mu)).$ (27)
Further, suppose $\eta$ does not charge the singleton $\{0\}$, so $P_{\mu,\sigma^{2}}(X_{i}=0)$ is always 00. The following was observed by Lai [1976, Section 5].
Lemma 4.4 (Density of the scale invariant reduction).
Let $V_{n}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the function
$V_{n}(x_{1},\dots,x_{n})=(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|).$ (28)
(It does not matter how $V_{n}$ is defined when $x_{1}=0$.)
Let $Q_{\mu,\sigma^{2}}^{n}$ be the push-forward measure of $P_{\mu,\sigma^{2}}^{\otimes n}$ under the map $V_{n}$. Let $\eta^{\langle n-1\rangle}$ be the measure on {±$\{\pm 1\}\times\mathbb{R}^{n-1}$ that charges both $\{1\}\times\mathbb{R}^{n-1}$ and $\{-1\}\times\mathbb{R}^{n-1}$ with $\eta^{\otimes(n-1)}$. Then,
$\frac{\mathrm{d}Q_{\mu,\sigma^{2}}^{n}}{\mathrm{d}\eta^{\langle n-1\rangle}}(x_{1},\dots,x_{n})=\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}/\tau-\mu/\sigma))\right\}\frac{\mathrm{d}\tau}{\tau}=\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{\tau\mu/\sigma,\tau^{2}}}{\mathrm{d}\eta}(x_{i})\right\}\frac{\mathrm{d}\tau}{\tau}.$ (29)
The lemma says, in probabilistic terms, that the “maximal invariant” reduction of the sample, $(X_{1}/|X_{1}|,X_{2}/|X_{1}|,\dots,X_{n}/|X_{1}|)$, has a density that relates to the density of the original sample in the form of (29) — which depends on $\mu$ and $\sigma^{2}$ only through $\mu/\sigma$.
Denoting the class of distributions having the same standardized mean by $\mathcal{P}=\{P_{\mu^{\prime},{\sigma^{\prime}}^{2}}:\mu^{\prime}/\sigma^{\prime}=\mu/\sigma\}$, we can further write the right hand side of (29) in the form of
$\int_{\mathcal{P}}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P}{\mathrm{d}\eta}(x_{i})\right\}\mathrm{J}(\mathrm{d}P).$ (30)
Where $\mathrm{J}$ is the Jeffreys prior over $\mathcal{P}$ with density $\frac{1}{\sigma(P)}$. This improper prior is known in the Bayesian literature as an uninformative prior on the scale parameter. We have shown above that taking the mixture (in some sense) with the Jeffreys prior $\int_{\tau>0}(\dots)\frac{\mathrm{d}\tau}{\tau}$ of likelihood ratios is equivalent to taking the likelihood ratio of the scale invariant reduction $X_{2}/|X_{1}|,\dots,X_{n}/|X_{1}|$. We shall make further remarks on how our approaches are related to previous Bayesian work with Jeffreys prior in Section 4.6.
Now using the fact that the general non-i.i.d. likelihood ratios are martingales (Lemma 2.5) on $Q_{\mu,\sigma^{2}}^{n}$, we have:
Lemma 4.5 (Scale invariant likelihood ratio).
For any $\theta$ and $\theta_{0}$, the process
$h_{n}(\theta;\theta_{0})=\frac{\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{\tau\theta,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}}{\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{\tau\theta_{0},\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}}$ (31)
is an NM for $\{P_{\mu_{0},\sigma_{0}^{2}}:\mu_{0}/\sigma_{0}=\theta_{0}\}$ on the scale invariant filtration $\{\mathcal{F}^{*}_{n}\}_{n\geqslant 1}$ with expected value 1.
4.4 An Extended Test Martingale for t-Test
It is now convenient for us to replace the general distribution $P_{\mu,\sigma^{2}}$ in the previous subsection with the Gaussian $N_{\mu,\sigma^{2}}$, and $\eta$ with the Lebesgue measure on $\mathbb{R}$. A direct calculation with the Gaussian density function in Lemma 4.5 gives the following,
Corollary 4.6 (Scale invariant t-likelihood ratio).
Let $\theta$ be any real number. The process $\{h_{\theta,n}\}_{n\geqslant 1}$, defined by
$h_{\theta,n}=\frac{\mathrm{e}^{-n\theta^{2}/2}}{\Gamma(n/2)}\int_{y>0}y^{n/2-1}\exp\left\{-y+\theta S_{n}\sqrt{\frac{2y}{V_{n}}}\right\}\mathrm{d}y,$ (32)
is an NM for $\mathcal{N}_{\mu=0}$ on the scale invariant filtration $\{\mathcal{F}^{*}_{n}\}_{n\geqslant 1}$.
The $\theta$ in the martingale (32) parametrizes the standardized mean $\mu/\sigma$ of the alternative. That is, when the actual distribution is in the class $\mathcal{N}_{\mu/\sigma=\theta}$, the process $\{h_{\theta,n}\}$ grows the fastest. A flat integral over $\theta$ yields an ENSM that stands behind Lai’s t-CS (Theorem 4.1), which we restate in our language as follows.
Theorem 4.7 (Scale invariant t-test extended martingale).
The process $\{H_{n}\}_{n\geqslant 1}$, defined as $H_{1}=\infty$ and
$H_{n}=\sqrt{\frac{2\pi}{n}}\left(\frac{nV_{n}}{nV_{n}-S_{n}^{2}}\right)^{n/2},\quad(\text{for }n\geqslant 2)$ (33)
is an ENSM for $\mathcal{N}_{\mu=0}$ on the scale invariant filtration $\{\mathcal{F}^{*}_{n}\}_{n\geqslant 1}$. Lai’s t-CS stated in Theorem 4.1 follows from applying the extended Ville’s inequality (Lemma 2.2) to this ENSM.
To see how $\{H_{n}\}$ behaves under $\mathcal{N}_{\mu=0}$, recall that $T_{n-1}=\sqrt{n-1}\frac{S_{n}}{\sqrt{nV_{n}-S_{n}^{2}}}$ has Student’s t-distribution of $n-1$ degrees of freedom. $H_{n}$ can thus be re-expressed as
$H_{n}=\sqrt{\frac{2\pi}{n}}\left(1+\frac{T_{n-1}^{2}}{n-1}\right)^{n/2}.$ (34)
$\{H_{n}\}$ works favorably as a test process for the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following symptotic result on both its growth index under the alternative and its convergence under the null.
Proposition 4.8 (Asymptotic behavior of the scale invariant t-test extended martingale).
Under any $N_{\mu,\sigma^{2}}$,
$\lim_{n\to\infty}\frac{\log H_{n}}{n}=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})\quad\text{almost surely}.$ (35)
Consequently, $\{H_{n}\}$ diverges almost surely to $H_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu\neq 0}$. Furthermore, $\{H_{n}\}$ converges almost surely to $H_{\infty}=0$ under $\mathcal{N}_{\mu=0}$.
The reader may compare the ENSM $\{H_{n}\}$ with the ENSM for the Z-test case by Wang and Ramdas [2023a, Proposition 5.7], obtained via a flat mixture over the standard Gaussian likelihood ratio. Both are free of any parameter. Under the null, both, as extended martingales, start from $\infty$ and shrink to 0 almost surely. Under the alternative, both start from $\infty$ and diverge back to $\infty$. Both can be seen as frequentist embodiments of the Bayesian idea of uninformative, improper prior.
We remark that the reduction of filtration from $\{\mathcal{F}_{n}\}$ to $\{\mathcal{F}^{*}_{n}\}$ does limit the stopping rules that are safe when working with the test processes. Stopping on a scale variant event would violate the safety of $\{H_{n}\}$. However, the effect of filtration reduction seems to disappear after applying extended Ville’s inequality to obtain a CS. This is because Ville’s inequality or extended Ville’s inequality, at a fundamental level, makes a claim only about the first exit time from $[0,\varepsilon]$ of the process (NSM or ENSM) — a stopping time on the process itself, one adapted to the canonical filtration of the process which is even coarser than $\{\mathcal{F}^{*}_{n}\}$.
4.5 Classical Test Martingales for t-Test
In some sense a classical, integrable test martingale issued at 1 is preferred. Besides a simple, universally valid rejection rule “reject when the test process exceeds $1/\alpha$”, the use of classical Ville’s inequality often leads to closed-form CSs as opposed the one in Theorem 4.7 that involves root finding. Further, these classical NMs often come with “tunable hyperparameters” arising from the mixture distributions. We replace the flat mixture on $\theta$ that leads to Theorem 4.7 by a Gaussian one, obtaining the following classical test martingales.
Theorem 4.9 (Scale invariant t-test martingale).
For any $c>0$, the process $\{G_{n}^{(c)}\}_{n\geqslant 1}$ defined by
$G_{n}^{(c)}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\frac{(n+c^{2})V_{n}}{(n+c^{2})V_{n}-S_{n}^{2}}\right)^{n/2}$ (36)
is an NM for $\mathcal{N}_{\mu=0}$ on the scale invariant filtration $\{\mathcal{F}^{*}_{n}\}_{n\geqslant 1}$ with expected value $\mathbb{E}[G_{n}^{(c)}]=G_{1}^{(c)}=1$. Consequently, let
$\operatorname{radius}_{n}=\sqrt{\frac{(n+c^{2})\left(1-\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}\right)}{\left\{\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}\right\}\vee 0}\left(\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)},$ (37)
the intervals
$\left[\widehat{\mu}_{n}\pm\operatorname{radius}_{n}\right]$ (38)
form a $(1-\alpha)$-CS for $\mu$ over $\mathcal{N}$. (When the denominator in (37) takes 0, $\operatorname{radius}_{n}=\infty$, the CI is the entire $\mathbb{R}$ at time $n$.)
While the test martingale (36) works without any limitation on $n$ or $c$, the CS (38) is non-trivial only when the $\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}$ term in the denominator is positive. If a CI at time $n$, or the width of the CS at some fixed time $n$ is of interest, one can take $c^{2}$ such that
$\frac{c^{2}}{n+c^{2}}=\alpha^{2/(n-1)}2^{-n/(n-1)},$ (39)
in which case
$\operatorname{radius}_{n}=\sqrt{\left(\alpha^{-2/(n-1)}2^{n/(n-1)}-2\right)\left(\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)},$ (40)
which is always finite and has a growth rate of $\alpha^{-1/(n-1)}$ as $\alpha\to 0$ while keeping the sample fixed, a slightly worse rate compared to the $\alpha^{-1/n}$ of the universal inference CS, and at least as good as the $\alpha^{-m/n(m-1)}$ of Lai’s original CS since $m\leqslant n$. On the other hand, if $c$ and $n$ are both fixed, the radius grows like $(\alpha-\alpha_{\star}(n))^{-1/n}$ as $\alpha\downarrow\alpha_{\star}(n)=\exp({\Theta}(-n\log n))$.
For a fixed $\alpha$ and a fixed $c$, the CS is non-trivial on $n\geqslant n_{0}$ for some $n_{0}$, instead of all $n\geqslant 1$. For example, when $c=0.01$ and $\alpha=0.05$, the range of (38) is $n\geqslant 3$. Thus, the starting time $m$ in Theorem 4.1 seems to be avoided by switching from the extended Ville’s inequality to the classical Ville’s inequality, but it remains in another form. In this case, if $n\to\infty$, $\operatorname{radius}_{n}$ shrinks as a function of $n$ at the rate of
$\sqrt{1-\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}}\approx\sqrt{\frac{\log(n/\alpha^{2}c^{2})}{n}}.$ (41)
One may further express $G_{n}^{(c)}$ in terms of the t-statistic under $\mathcal{N}_{\mu=0}$ as well,
$G_{n}^{(c)}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1+\frac{n}{\frac{(n+c^{2})(n-1)}{T_{n-1}^{2}}+{c^{2}}}\right)^{n/2}$ (42)
which again leads to the following asymptotic result that shows their merits for being test process candidates.
Proposition 4.10 (Asymptotic behavior of the scale invariant t-test martingales).
Under any $N_{\mu,\sigma^{2}}$,
$\lim_{n\to\infty}\frac{\log G_{n}^{(c)}}{n}=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})\quad\text{almost surely}.$ (43)
Consequently, $\{G_{n}^{(c)}\}$ diverges almost surely to $G^{(c)}_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu\neq 0}$. Furthermore, $\{G_{n}^{(c)}\}$ converges almost surely to $G^{(c)}_{\infty}=0$ under $\mathcal{N}_{\mu=0}$.
The NM $\{G_{n}^{(c)}\}$ thus have the same asymptotic properties as the ENSM $\{H_{n}\}$, but the free parameter $c$, which is absence for the ENSM, does introduce a difference that emerges only non-asymptotically. To wit, if one fixes $n$ and the data-dependent quantity $T^{2}_{n-1}$, the value of $G_{n}^{(c)}$ would approach 0 (hence the power vanishes) if $c$ is too large or too small.
The reader may compare Theorem 4.9 with Theorem 4.7, and compare this comparison with the comparison between the Gaussian mixed NSM [Wang and Ramdas, 2023a, Proposition 5.6] and flat mixed ENSM [Wang and Ramdas, 2023a, Proposition 5.7] in the Z-test case. Multiple similarities manifest. A full comparison shall be presented next in Table 3. It is unclear why Lai [1976] skipped this more universally accepted method of a proper Gaussian mixture and used an improper flat mixture instead, which seems very ahead of its time in hindsight.
We also note that Lindon et al. [2022, Proposition 3.7] concurrently derive the same t-test martingale and confidence sequence as our Theorem 4.9, a special case of their F-test martingale. Our paper, however, is concerned much more with comparing the Gaussian mixture with Lai’s flat mixture in theory and practice, with understanding the tradeoffs between test martingales in the reduced filtration and e-processes in the original filtration for this fundamental problem, as well as deriving their (often suboptimal) dependence on $\alpha$.
In place of the full Gaussian prior, a half-Gaussian prior mixture will give us a “semi-one-sided” test process in the sense that it is valid only for the point null $\mathcal{N}_{\mu=0}$ (unlike Theorem 3.4) but still leads to a one-sided confidence sequence.
Theorem 4.11 (Scale invariant semi-one-sided t-test e-process).
For any $c>0$, the process $G_{n}^{(c-)}$ defined by
$G_{n}^{(c-)}=2\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}-\left(1-\frac{(S_{n}\wedge 0)^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}\right)=2G_{n}^{(c)}-2G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}.$ (44)
is an e-process for $\mathcal{N}_{\mu=0}$ on the scale invariant filtration $\{\mathcal{F}_{n}^{*}\}_{n\geqslant 1}$.
Here, the notation $G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}$ refers to replacing the sample sum $S_{n}$ in the two-sided martingale (36) with its negative part $S_{n}\wedge 0$. The process is an e-process instead of a martingale because a step of approximation is used to unravel a non-closed-form martingale, as seen from its proof in Section C.1. As a direct consequence of Proposition 4.10, the process $G_{n}^{(c-)}=2G_{n}^{(c)}-2G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}$ diverges almost surely to $G^{(c-)}_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu>0}$ and converges almost surely to $G^{(c-)}_{\infty}=0$ under $\mathcal{N}_{\mu=0}$, and satisfies exactly the same asymptotic as Proposition 3.5,
$\lim_{n\to\infty}\frac{\log G^{(c-)}_{n}}{n}=\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})\quad\text{almost surely}.$ (45)
4.6 Remark: Bayesian t-Test with the JZS Prior, and Cauchy Mixture
In a highly influential paper, Rouder et al. [2009] provide a Bayesian framework for the t-test that makes extensive use of Bayes factors. To explain this approach, we consider as we did in Section 4.3 a general location-scale family $P_{\mu,\sigma^{2}}$ dominated by a reference measure $\eta$. The Bayes factor for the null $\mathcal{P}_{\mu=0}=\{P_{0,\sigma^{2}}:\sigma>0\}$ and the alternative $\mathcal{P}_{\mu\neq 0}$ is defined as [Rouder et al., 2009, p.229]
$B_{n}=\frac{M^{0}_{n}}{M^{1}_{n}}=\frac{\int\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{0,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\pi_{0}(\mathrm{d}\tau^{2})}{\int\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{m,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\pi_{1}(\mathrm{d}m,\mathrm{d}\tau^{2})},$ (46)
where $\pi_{0}$ and $\pi_{1}$ are priors on $\mathbb{R}^{+}$ and $\mathbb{R}\times\mathbb{R}^{+}$, that can be chosen freely by the statistician. Rouder et al. [2009, p.231], regarding themselves as “objective Bayesians”, recommend the following choice:
$\pi_{0}(\mathrm{d}\tau^{2})=\tau^{-2}\mathrm{d}\tau^{2}=2\tau^{-1}\mathrm{d}\tau,\quad\text{and}\quad\pi_{1}(\mathrm{d}m,\mathrm{d}\tau^{2})=C_{0,1}(\mathrm{d}(m/\tau))\cdot\pi_{0}(\mathrm{d}\tau^{2}),$ (47)
where $C_{0,1}$ is the standard Cauchy distribution. This is dubbed the “JZS prior”, an acronym for the Jeffreys prior $\pi_{0}$, and the Cauchy prior on $\mu/\sigma$ due to Zellner and Siow [1980]. We can immediately write $B_{n}$ as
$B_{n}^{\text{JZS}}=\frac{\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{0,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}}{\int_{\theta}\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{\tau\theta,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}C_{0,1}(\mathrm{d}\theta)}.$ (48)
Comparing to mixing the scale-invariant likelihood ratio (31) (letting $\theta_{0}=0$) with a prior $\varpi$ on the alternative $\theta$,
$M_{n}^{\varpi}=\int h_{n}(\theta;0)\varpi(\mathrm{d}\theta)=\frac{\int_{\theta}\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{\tau\theta,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}\varpi(\mathrm{d}\theta)}{\int_{\tau>0}\left\{\prod_{i=1}^{n}\frac{\mathrm{d}P_{0,\tau^{2}}}{\mathrm{d}\eta}(X_{i})\right\}\frac{\mathrm{d}\tau}{\tau}}.$ (49)
Clearly, $B_{n}^{\text{JZS}}$ equals $1/M_{n}^{\varpi}$ when the mixture measure $\varpi$ is taken to be $C_{0,1}$. Some remarks regarding the comparison. First, Rouder et al. [2009] did not mention the sequential benefits of their approach, as $\{1/B_{n}^{\text{JZS}}\}$ is an NM for $\mathcal{P}_{\mu=0}$ by virtue of Lemma 4.5, or equivalently $\{B_{n}^{\text{JZS}}\}$ is an anytime-valid p-value which we briefly defined in Section 2.2. Second, while choosing $\pi_{0}$ to be the Jeffreys prior seems necessary to attain a test process according to our Section 4.3, different (perhaps objectivistic) methodological choices lead to different priors on $\theta=\mu/\sigma$. Lai’s choice in Section 4.4 is a flat $\varpi$ while in Section 4.5 we choose $\varpi=N_{0,c^{-2}}$, both leading to closed-form expressions. Rouder et al.’s Cauchy prior, in turn, arises from a hyper-prior on the precision $c^{-2}$ of this Gaussian prior $\theta\sim N_{0,c^{-2}}$, which is $c^{2}\sim\chi^{2}_{1}$ proposed by Zellner and Siow [1980], but this has led to the $B_{n}^{\text{JZS}}$ that lacks a closed form.
5 Comparison of Results
5.1 Theoretical Comparison
We have presented three t-confidence sequences (i.e., confidence sequences for $\mu$ over $\mathcal{N}$) so far, based on universal inference and scale invariance, and we summarize them in Table 1. Two additional results that can be obtained from (1) the two-step plug-in method of Jourdan et al. [2023, Appendix E], (2) the confidence sequence for median by Howard and Ramdas [2022, Theorem 1] are also included, both described in Appendix A. In terms of shrinkage rates, these two results enjoy the optimal iterated logarithmic rate faster than the other three. However, we point out that this is achieved by the technique of “stitching” which universally turns a $n^{-1/2}$ CI into $n^{-1/2}\log\log(n)$ CS (by paying a significant price of constants), thus applicable to the other three methods as well, see e.g. Howard et al. [2021, Section 3.1] and Duchi and Haque [2024, Section 4]. Growth-wise, the original flat-mixed CS by Lai [1976] and the result of Jourdan et al. [2023] are subject to a parameter-dependent worse rate, and the result of Howard and Ramdas [2022] becomes degenerate when $\alpha$ is small enough, thus they are less favorable when $\alpha$ is very small (an issue that Theorem 4.9 can avoid by tuning $c^{2}$.).
Result Theorem 3.2 Theorem 4.1 [Lai, 1976] Theorem 4.9 Theorem A.3 [Jourdan et al., 2023] Proposition A.4 [Howard and Ramdas, 2022]
Method Universal inference Scale invariant likelihood mixture Likelihood mixture (mean) $\chi^{2}$ Chernoff bound (variance) Chernoff bound for median
Mixture N/A Flat Gaussian N/A
Free parameters Point estimators $\tilde{\mu}_{i}$ and $\tilde{\sigma}_{i}^{2}$ Starting time $m\geqslant 2$ Prior precision $c^{2}$ Stitching parameters $\eta>0,s>1$ N/A
Growth rate ($\alpha\to 0$, or $\alpha\to\alpha_{\star}$ $=\mathrm{e}^{-\widetilde{\Theta}(n)}$ while $n$ fixed) $\alpha^{-1/n}$ $\alpha^{-m/n(m-1)}$ $\alpha^{-1/(n-1)}$ (CI) $(\alpha-\alpha_{\star})^{-1/n}$ (CS) $\alpha^{-(1+\eta)/(n-1)}$ $\sqrt{\log(1/(\alpha-\alpha_{\star}))}$
Shrinkage rate ($n\to\infty$ while $\alpha$ fixed) $n^{-1/2}\operatorname*{polylog}(n)$ $n^{-1/2}\log\log(n)$
Table 1: Comparison of t-confidence sequences in this paper.
We now zoom out and cross-compare sequential t-tests with sequential Z-tests in terms of their test processes. As we mentioned earlier, both universal inference and likelihood mixture can be used on Z-test and t-test to construct e-processes or extended e-processes, and confidence sequences. Additional results on universal inference Z-tests may be found in Appendix B. We summarize in Tables 2 and 3 the results from universal inference and likelihood mixture respectively. Note that the plug-in method of Jourdan et al. [2023] and the median bound of Howard and Ramdas [2022] yield confidence sequences without any test process, so are not included in these tables.
Result Corollary 3.1 Proposition B.2 Theorem 3.2 Theorem 3.4
Problem Sequential Z-test Sequential t-test
Null $N_{0,\sigma^{2}}$ $\mathcal{N}_{\mu\leqslant 0,\sigma^{2}}$ $\mathcal{N}_{\mu=0}$ $\mathcal{N}_{\mu\leqslant 0}$
Alternative $\mathcal{N}_{\mu\neq 0,\sigma^{2}}$ $\mathcal{N}_{\mu>0,\sigma^{2}}$ $\mathcal{N}_{\mu\neq 0}$ $\mathcal{N}_{\mu>0}$
Test process; class; filtration $\sigma^{n}\mathrm{e}^{\frac{V_{n}}{2\sigma^{2}}}K_{n}$ NM $\{\mathcal{F}_{n}\}$ $\sigma^{n}\mathrm{e}^{\frac{V_{n}-n(\widehat{\mu}_{n}\wedge 0)^{2}}{2\sigma}}K_{n}$ e-process $\{\mathcal{F}_{n}\}$ $\left(\overline{X^{2}_{n}}\right)^{n/2}\mathrm{e}^{n/2}K_{n}$ e-process $\{\mathcal{F}_{n}\}$ $\left(\overline{X^{2}_{n}}-(\widehat{\mu}_{n}\wedge 0)^{2}\right)^{n/2}\mathrm{e}^{n/2}K_{n}$ e-process $\{\mathcal{F}_{n}\}$
Behavior under null at most $1\to 1/\alpha$ with probability $1-\alpha$
Behavior under alternative $1\to\infty$ a.s.
Growth index $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
Table 2: Comparison of universal inference-based Z-tests and t-tests, where $K_{n}:=\prod_{i=1}^{n}\frac{\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\widetilde{\sigma}_{i-1}}$. The “growth index” of a process $M_{n}$ refers to the almost sure limit of $\frac{\log M_{n}}{n}$ under the true distribution $N_{\mu,\sigma^{2}}$.
Result WR23 Prop. 5.6 WR23 Prop. 5.7 WR23 Prop. 6.3 Theorem 4.9 Theorem 4.7 Theorem 4.11
Problem Sequential Z-test Sequential t-test
Null $N_{0,\sigma^{2}}$ $\mathcal{N}_{\mu\leqslant 0,\sigma^{2}}$ $\mathcal{N}_{\mu=0}$
Alternative $\mathcal{N}_{\mu\neq 0,\sigma^{2}}$ $\mathcal{N}_{\mu>0,\sigma^{2}}$ $\mathcal{N}_{\mu\neq 0}$ $\mathcal{N}_{\mu>0}$
Base martingales; parametrized by Gaussian likelihood ratio; $\mu$, the alternative mean Scale invariant likelihood ratio; $\theta=\mu/\sigma$, the alternative standardized mean
Mixture $\mu\sim N_{0,c^{-2}}$ $\mu\sim F$ $\mu\sim F^{>0}$ $\theta\sim N_{0,c^{-2}}$ $\theta\sim F$ $\theta\sim N_{0,c^{-2}}^{>0}$
Test process; class; filtration $\sqrt{\frac{c^{2}}{n+c^{2}}}\mathrm{e}^{\frac{n^{2}\widehat{\mu}_{n}^{2}}{2(n+c^{2})\sigma^{2}}}$ NM $\{\mathcal{F}_{n}\}$ $\frac{1}{\sqrt{n}}\mathrm{e}^{\frac{n\widehat{\mu}_{n}^{2}}{2\sigma^{2}}}$ ENSM $\{\mathcal{F}_{n}\}$ $\frac{1}{\sqrt{4n}}V(\sqrt{n}\widehat{\mu}_{n}/\sigma)$ ENSM $\{\mathcal{F}_{n}\}$ $\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\frac{(n+c^{2})V_{n}}{(n+c^{2})V_{n}-S_{n}^{2}}\right)^{\frac{n}{2}}$ NM $\{\mathcal{F}_{n}^{*}\}$ $\sqrt{\frac{2\pi}{n}}\left(\frac{nV_{n}}{nV_{n}-S_{n}^{2}}\right)^{\frac{n}{2}}$ ENSM $\{\mathcal{F}_{n}^{*}\}$ $2G_{n}^{(c)}-2G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}$ e-process $\{\mathcal{F}_{n}^{*}\}$
Behavior under null $1\to 0$ a.s. $\infty\to 0$ a.s. $1\to 0$ a.s. $\infty\to 0$ a.s. $1\to 0$ a.s.
[…]
Growth index $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
Table 3: Comparison of mixture-based Z-tests and t-tests. WR23 refers to the paper by Wang and Ramdas [2023a], which also covers the subGaussian case, but we omit it for a more direct comparison. The function $V$ is $V(x)=\mathrm{e}^{x^{2}/2}(1+\operatorname*{erf}(x/\sqrt{2}))$, and $G_{n}^{(c)}$ refers to the test process listed above under Theorem 4.9.
5.2 Simulations
Figure 1: Five classes of confidence sequences for t-test under $N_{0,1}$ observations.
Refer to captionRefer to caption
Figure 2: This pair of plots studies the behavior of width as one of $\alpha$ and $n$ vary, holding the other fixed. In the left plot, we plot widths of 3 CSs against $\alpha$ at $n=500$; whereas in the right plot, we plot widths multiplied by $\sqrt{n}$ against sample size $n$ at $\alpha=0.05$. Both are under $N_{0,1}$ observations, repeated 25 times. Lines indicate averages over 25 repeats.
We demonstrate the three confidence sequences in the paper, as well as the additional “plug-in” one, Theorem A.3, and the median CS, Proposition A.4, by some experiments. For the universal inference CS, we calculate the point estimators by putting a prior where we “imagine” having seen pre-observations $-1$ and $1$, and defining $\tilde{\mu}_{i}$ and $\tilde{\sigma}_{i}^{2}$ as the empirical means and variances posterior. For the Gaussian mixture of scale invariant likelihood ratios, we take $c^{2}$ to be 1 and 100.
First, we present them visually in the plot of Figure 1 in a single run of i.i.d. standard normal observations. For the plug-in CS, we fix $\eta=0.5$ and let $s\in\{1.25,4\}$ since $s$ is the parameter that controls the shrinkage rate. We see that the Gaussian mixture CS we derived in Section 4.5 with $c=1$ performs better than Lai’s improper, flat mixture one; increasing $c$ results in interval explosion at earlier times but is slightly tighter at later times. The universal inference CS seems unfavorable at later times, but we shall soon see in repeated experiments that this is not necessarily the case. The plug-in CS by Jourdan et al. [2023], however, is much looser than the rest even at $n=10000$ despite its asymptotically optimal iterated logarithmic rate, thus is excluded from the next simulation. A similar issue of large constant is observed in the iterated-logarithmic-rate median CS by Howard and Ramdas [2022] as well.
We next compare their rates of growth and shrinkage. Still drawing observations from i.i.d. $N_{0,1}$, we first fix $n=500$ and let $\alpha$ vary, then fix $\alpha=0.05$ and let $n$ vary, plotting the widths of the CSs in 20 independent runs, since all of these CSs have random widths. The results are shown in Figure 2. We see with greater clarity that Lai’s CS is looser than Gaussian mixture-based CSs especially when $\alpha$ is small, even though the extended Ville’s inequality seems a tighter and more advanced technique; larger precision $c^{2}$ in Gaussian mixture gradually gains an advantage as $n$ increases. Even after a “smoothing” prior, the universal inference CS is still the most volatile. Its width varies significantly among runs.
6 Optimality
In this section, we show that the $\mathcal{O}(\alpha^{-1/n})$ or $\mathcal{O}(\alpha^{-1/(n-1)})$ growth rate (when $n$ is fixed and $\alpha\to 0$) in our t-confidence sequences shown in Table 1, and that the $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ growth index in our t-test processes shown in Tables 2 and 3, are both optimal. These two kinds of optimality require different techniques, but the Kullback-Leibler divergence
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\mu^{2}+\sigma^{2}})=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ (50)
is used in both cases.
6.1 Information-Theoretic Lower Bound of t-Confidence Intervals
Recall from Table 1 that our best CSs have width scaling with $\alpha$ like the nonstandard rate $\alpha^{-1/n}$, which is very different from the rate of $\sqrt{\log(1/\alpha)}$ that one typically sees for confidence intervals and sequences when the variance is known. Indeed, we will see now that for t-CIs (and hence t-CSs), this $\alpha^{-1/n}$ rate is actually minimax optimal.
We define $\mathcal{L}_{\alpha,n}$ to be the set of all lower $(1-\alpha)$-confidence bounds for $\mu$ over $\mathcal{N}$ with fixed sample size $n$, i.e., functions $L:\mathbb{R}^{n}\to\mathbb{R}$ such that
for $\text{for all }P\in\mathcal{N},\quad P(\mu(P)\geqslant L(X_{1},\dots,X_{n}))\geqslant 1-\alpha.$ (51)
Fix a particular $\varepsilon>0$ and $P\in\mathcal{N}$. For any $L\in\mathcal{L}_{\alpha,n}$, denote by $W_{\varepsilon}(L,P)$ the $\varepsilon$-upper quantile of the random variable $\mu(P)-L(X_{1},\dots,X_{n})$ under $P$, i.e.,
$P(\mu(P)-L(X_{1},\dots,X_{n})\geqslant W_{\varepsilon}(L,P))=\varepsilon.$ (52)
Let us study the minimax rate
$M^{-}_{\alpha,n}=\inf_{L\in\mathcal{L}_{\alpha,n}}\sup_{P\in\mathcal{N}}\sigma^{-1}(P)\cdot W_{2\alpha}(L,P).$ (53)
The definition above characterizes the following: the event that a lower CI covers $\mu$ with an excess of $M^{-}_{\alpha,n}$ has probability at least $2\alpha$, which is an event disjoint from the event that it fails to cover $\mu$ which has probability at most $\alpha$. As we shall soon see, this one-sided lower bound immediately implies a lower bound for two-sided CIs. The two layers of $2\alpha$-low probability and $(1-\alpha)$-high probability are introduced to include confidence intervals with data-dependent radii, which, though uncommon in Z-tests, recur in t-tests throughout this paper.
We establish the following lower bound for $M^{-}_{\alpha,n}$:
Theorem 6.1 (Minimax Lower Bound of Lower t-CIs).
If $\alpha<1/3$, $M^{-}_{\alpha,n}\geqslant\sqrt{(6\alpha-9\alpha^{2})^{-2/n}-1}$.
This lower bound is proved by employing a standard two-point information-theoretic argument, via a pair of $N_{\mu,\sigma^{2}}$ and $N_{0,\mu^{2}+\sigma^{2}}$ such that the total variation distance between $N_{\mu,\sigma^{2}}^{\otimes n}$ and $N_{0,\mu^{2}+\sigma^{2}}^{\otimes n}$ is at most $1-3\alpha$. The details are in Section C.4.
Immediately following the one-sided lower bound, we can establish the following two-sided lower bound. Let $\mathcal{I}_{\alpha,n}$ be the set of all two-sided $(1-\alpha)$-confidence intervals for $\mu$ over $\mathcal{N}$ with fixed sample size $n$. For any $\operatorname{CI}\in\mathcal{I}_{\alpha,n}$, denote by $W_{\varepsilon}(\operatorname{CI},P)$ the $\varepsilon$-upper quantile of its width $|\operatorname{CI}|$ under $P$, and let
$M_{\alpha,n}=\inf_{\operatorname{CI}\in\mathcal{I}_{\alpha,n}}\sup_{P\in\mathcal{N}}\sigma^{-1}(P)\cdot W_{\alpha}(\operatorname{CI},P).$ (54)
Then, we have the following minimax lower bound, also proved in Section C.4:
Corollary 6.2 (Minimax Lower Bound of t-CIs).
If $\alpha<1/3$, $M_{\alpha,n}\geqslant\sqrt{(6\alpha-9\alpha^{2})^{-2/n}-1}$.
If $n$ is fixed, the lower bound above is asymptotically $\mathcal{O}(\alpha^{-1/n})$, the growth rate attained by universal inference method (Theorem 3.2), and almost attained by Gaussian mixture of scale invariant likelihood ratios method (Theorem 4.9) in our paper.
6.2 The Surprising $\alpha$-Suboptimality of the Classical t-Test
While the lower bound $\mathcal{O}(\alpha^{-1/n})$ proved above is matched by one of the t-CSs we derive, we now come to the surprising fact that the confidence interval obtained by inverting the classical one-sample Student’s t-test with a fixed sample size $n$ fails to attain the same optimality as a function of $\alpha$. Under $N_{\mu,\sigma^{2}}$, the classical t-test relies on the following “pivotal” quantity that follows the t-distribution of $n-1$ degrees of freedom:
$T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu}{\sqrt{nV_{n}-S_{n}^{2}}},$ (55)
where $S_{n}$ and $V_{n}$ denote the sum of samples and the sum of sample squares respectively. The following, therefore, is a $(1-\alpha)$-confidence interval for $\mu$:
$\operatorname{CI}_{n}^{\mathsf{t}}=\left[\widehat{\mu}_{n}\pm\sqrt{\frac{nV_{n}-S_{n}^{2}}{n^{2}(n-1)}}\cdot t_{1-\alpha/2}(n-1)\right],$ (56)
where $t_{1-\alpha/2}(n-1)$ denotes the $21-\alpha/$ 2 quantile of the student t-distribution with $n-1$ degrees of freedom. Since the survival function of the student t-distribution with $n-1$ degrees of freedom scales as $\mathcal{O}(x^{-(n-1)})$, $t_{1-\alpha/2}(n-1)$, therefore the width of $\operatorname{CI}_{n}^{\mathsf{t}}$ as well, scales as $\mathcal{O}(\alpha^{-1/(n-1)})$ when the sample is fixed and $\alpha\to 0$.
The $\alpha^{-1/(n-1)}$ growth rate of the classical t-CI is the same as the Gaussian mixture CS, Theorem 4.9, but worse than the $\alpha^{-1/n}$ optimal rate attained by the other of our t-CSs, i.e., Theorem 3.2 by universal inference. That is, if we fix a sample size $n$ and decrease $\alpha$, the CI reported by Theorem 3.2 will eventually be tighter than $\operatorname{CI}_{n}^{\mathsf{t}}$ when $\alpha$ is sufficiently small, as is shown in Figure 3.
Figure 3: Growth rates of the t-CS by universal inference (Theorem 3.2) and the classical t-test CI with $n=3$ observations.
We explain this phenomenon from the following perspective. Consider the case with only $n=1$ observation $X_{1}$. The classical t-test is undefined; however, Theorem 3.2 implies the following valid e-value (taking $\widetilde{\mu}_{0}=0$ and $\widetilde{\sigma}_{0}=1$) for $N_{\mu,\sigma^{2}}$,
$|X_{1}-\mu|\exp\left(\frac{1-X_{1}^{2}}{2}\right).$ (57)
To see that this is an e-value without invoking the full Theorem 3.2, use the fact that $|x-\mu|\leqslant\sigma\exp((x-\mu)^{2}/2\sigma^{2}+1/2)$ for all $x,\sigma$ to conclude that
$\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(|X_{1}-\mu|\cdot\exp\left(\frac{1-X_{1}^{2}}{2}\right)\right)\leqslant$ (58)
$\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\sigma\exp\left(\frac{(X_{1}-\mu)^{2}}{2\sigma^{2}}+\frac{1}{2}\right)\cdot\exp\left(\frac{1-X_{1}^{2}}{2}\right)\right)=\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu,\sigma^{2}}}(X_{1})\right)=1.$ (59)
Applying Markov’s inequality to the e-value above gives rise to a non-trivial CI for $\mu$:
$\left[X_{1}\pm\alpha^{-1}\exp\left(\frac{X_{1}^{2}-1}{2}\right)\right]$ (60)
with $\mathcal{O}(\alpha^{-1})$ growth rate. It takes a second observation for the classical t-test to be non-trivial, thresholding the t-statistic $T_{1}=\frac{(X_{1}-\mu)+(X_{2}-\mu)}{|X_{1}-X_{2}|}$, which has a Cauchy distribution, resulting in a $\mathcal{O}(\alpha^{-1})$-CI that corresponds to the $\mathcal{O}(x^{-1})$ tail of Cauchy distribution. While with $n=2$, Theorem 3.2 produces a better CI in terms of growth rate: taking just $\widetilde{\mu}_{0}=0,\widetilde{\mu}_{1}=X_{1}$ and $\widetilde{\sigma}_{0}=\widetilde{\sigma}_{1}=1$ for simplicity, the e-value reads,
$\mathrm{E}_{N_{\mu,\sigma^{2}}}\left[\frac{(X_{1}-\mu)^{2}+(X_{2}-\mu)^{2}}{2}\exp\left(1-\frac{X_{1}^{2}+(X_{2}-X_{1})^{2}}{2}\right)\right]\leqslant 1.$ (61)
Applying Markov’s inequality to the e-value above gives rise to a CI with $\mathcal{O}(\alpha^{-1/2})$ growth rate due to the quadratic dependence on $\mu$.
6.3 Reverse Information Projection and Numeraire
We now show that the recurring growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ in various test processes presented thus far is not a coincidence, as it is the optimal rate of growth in the problem of t-test.
It is known that when testing the point null $P$, if the real distribution $Q\neq P$, an e-process $\{M_{n}\}$ grows optimally at the rate of
$\frac{\mathrm{E}_{Q}(\log M_{n})}{n}\leqslant D_{\operatorname{KL}}(Q\|P),$ (62)
with equality if and only if $\{M_{n}\}$ is the likelihood ratio in Lemma 2.3. See e.g. Vovk and Wang [2022, Lemma 2.1]. For example, in sequential Z-tests, $P=N_{0,\sigma^{2}}$ and $Q=N_{\mu,\sigma^{2}}$, then the upper bound reads
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}})=\frac{\mu^{2}}{2\sigma^{2}}.$ (63)
More generally, if we are to test a composite null $\mathcal{P}$ and the real distribution $Q\notin\mathcal{P}$, an e-process $\{M_{n}\}$ grows at most at the rate of
$\frac{\mathrm{E}_{Q}(\log M_{n})}{n}\leqslant\inf_{P\in\mathcal{P}}D_{\operatorname{KL}}(Q\|P).$ (64)
If the infimum $\inf_{P\in\mathcal{P}}D_{\operatorname{KL}}(Q\|P)$ is attained as minimum by some $P^{*}\in\mathcal{P}$, we say that $P^{*}$ is the reverse information projection of $Q$ onto the set $\mathcal{P}$; if, further, the likelihood ratio process $M_{n}^{*}=\prod_{i=1}^{n}\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{i})$ is an e-process for $\mathcal{P}$, then $\{M_{n}^{*}\}$ is the optimal (fastest growing) e-process for $\mathcal{P}$ under the real distribution $Q$, the one that attains equality in (64), and is called the numeriare e-process. This is due to the following result by Larsson et al. [2024]:
Lemma 6.3 (Theorem 4.1 of Larsson et al. [2024]).
If $Y=\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})$ is an e-value for $\mathcal{P}$ (i.e. $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(Y)\leqslant 1$) and $P^{*}\in\mathcal{P}$, then, among all e-values for $\mathcal{P}$, $Y$ maximizes $\mathrm{E}_{Q}(\log Y)$ and is called the numeraire e-value.
We have the following statement regarding the reverse information projection and the numeraire in t-tests where $\mathcal{P}=\mathcal{N}_{\mu=0}$ and $Q=N_{\mu,\sigma^{2}}$.
Proposition 6.4.
$P^{*}=N_{0,\sigma^{2}+\mu^{2}}$ attains the following minimum
$\min_{P\in\mathcal{N}_{\mu=0}}D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|P)=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}).$ (65)
thus it is the reverse information projection of $N_{\mu,\sigma^{2}}$ onto $\mathcal{N}_{\mu=0}$. Further, $M_{n}^{*}=\prod_{i=1}^{n}\frac{\mathrm{d}N_{\mu,\sigma^{2}}}{\mathrm{d}N_{0,\sigma^{2}+\mu^{2}}}(X_{i})$ is a supermartingale for $\mathcal{N}_{\mu=0}$, thus the numeraire e-process. Consequently, any e-process $\{M_{n}\}$ for $\mathcal{N}_{\mu=0}$ grows at most
$\frac{\mathrm{E}_{N_{\mu,\sigma^{2}}}(\log M_{n})}{n}\leqslant\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}),$ (66)
equality if and only if $\{M_{n}\}$ is the numeraire $\{M_{n}^{*}\}$.
The proposition above is proved by computing explicitly that $\mathrm{E}_{P_{0}}\left(\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})\right)\leqslant 1$, detailed in Section C.4. It indicates that all test processes summarized in Tables 2 and 3 have optimal growth indices.
7 Concluding Remarks
In this paper, we derive several sequential t-tests via universal inference and scale invariant likelihood ratio mixtures. In particular, we focus on the construction of test processes (nonnegative supermartingales, e-processes, and extended nonnegative supermartingales), and the confidence sequences they imply.
The two methods we explore yield similar results in numerous regards. They asymptotically share the same optimalities in terms of test processes and confidence sequences alike; they both extend to the one-sided test; and the computation of the test processes only requires linear time in both cases (unlike the quadratic cost that universal inference incurs worst-case, as mentioned by Wasserman et al. [2020, Section 7]).
One of the key differences between the two methods is the filtrations to which the processes are adapted, as the test processes scale invariant likelihood ratio mixtures produce are only adapted to a reduced filtration, allowing for a smaller set of stopping times that optionally stop the experiment. On the other hand, the universal inference test processes, while robust under a larger set of stopping times, sometimes suffer from lower stability depending on the quality of the estimators plugged in. We remark that the issue of filtration disappears once we convert tests into confidence sequences.
While we only focus on the class of one-dimensional normal distributions $\mathcal{N}$ in this paper, we expect similar techniques may lead to sequential Hotelling’s T-tests (with test processes and confidence sphere sequences [Chugg et al., 2023]) for the class of multivariate normal distributions. A slightly different way to generalize in dimensionality is the sequential F-tests for linear models, on which the most recent and comprehensive work by Lindon et al. [2022] only covers one of the numerous methods in our paper (namely, Gaussian mixture of scale invariant likelihood ratios). For potential nonparametric extensions of our work beyond normal distributions, one may need to appeal to asymptotic sequential tests and confidence sequences, proposed recently by Waudby-Smith et al. [2021] — whose Z-test-like results, however, arguably stand as nonparametric, asymptotic extensions of our t-tests already. We leave these topics for future authors to study in depth.
Acknowledgements
The authors acknowledge support from NSF grants IIS-2229881 and DMS-2310718. AR thanks Ashwin Pananjady for useful conversations.
Appendix A Additional Sequential t-Tests in the Literature
A.1 A Plug-in t-Confidence Sequence by Jourdan et al. [2023]
In this section, we review the method by Jourdan et al. [2023, Appendix E], who first derive a confidence sequence for the variance $\sigma^{2}$ based on the sample variance $v_{n}$ over $\mathcal{N}$ (using an example by Howard et al. [2021]), and then use a standard Z-confidence sequence via the method of likelihood mixture (e.g. Equation 10 of Howard et al. [2021]) that requires knowing the true variance $\sigma^{2}$. They then use a union bound to combine the two to obtain a rectangular confidence sequence for the pair $(\mu,\sigma^{2})$ over $\mathcal{N}$. A tighter confidence sequence can be obtained, we remark, if only the estimation of $\mu$ is of interest.
Throughout the section, we follow the same notation as Jourdan et al. [2023] of $\overline{W}_{i}(x)=-W_{i}(-\exp(-x))$ where $W_{0}$, $W_{-1}$ are the Lambert $W$ functions, the inverses of $x\mapsto x\mathrm{e}^{x}$. It is not hard to see that
$\lim_{x\to\infty}\frac{\overline{W}_{-1}(x)}{x}=1,\quad\lim_{x\to\infty}\frac{1/\overline{W}_{0}(x)}{\exp(x)}=1.$ (67)
We also let $\zeta$ be the Riemann Zeta function, which naturally arises in stitching [Howard et al., 2021], a technique for obtaining law-of-the-iterated-logarithm rate confidence sequences. Recall that we denote by $v_{n}$ the sample variance without Bessel correction,
$v_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}.$ (68)
Corollary 26 of Jourdan et al. [2023] states the following bound:
Proposition A.1 (Stitched Upper Confidence Sequence for Gaussian Variance).
Let $s>1,\eta>0$. The intervals
$\left[0,\frac{v_{n}}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}\right]$ (69)
with $n>n_{\min}$ where $n_{\min}=\operatorname*{polylog}(1/\alpha)$ is the largest $n$ such that the denominator above is nonpositive, form an $(1-\alpha)$-CS for $\sigma^{2}$ over $\mathcal{N}$.
It is very worth noting that, since $1/\overline{W}_{0}(x)\approx\exp(x)$ for large $x$, the confidence sequence (71) has polynomial, as opposed to the seemingly logarithmic dependence on $1/\alpha$.
For the mean $\mu$, Lemma 28 of Jourdan et al. [2023] quotes the following well-known bound:
Proposition A.2 (Stitched Normal Mixture Confidence Sequence for Gaussian Mean).
Let $s>1$. The intervals
$\left[\widehat{\mu}_{n}\pm{\sigma}\sqrt{\frac{\overline{W}_{-1}\left(1+2\log(1/\alpha)+2\log\zeta(s)+2s(1-\log(2s))+2s\log(2s+\log n)\right)}{n}}\right]$ (70)
form an $(1-\alpha)$-CS for $\mu$ over $\mathcal{N}_{\sigma^{2}}$.
A union bound between Propositions A.1 and A.2 immediately implies the following t-confidence sequence, by “plugging in” the variance upper confidence bound into the fixed-variance confidence sequence.
Theorem A.3 (Plug-In t-Confidence Sequence).
Let $s>1,\eta>0$. The intervals
$\left[\widehat{\mu}_{n}\pm\sqrt{\frac{v_{n}}{n}\cdot\frac{\overline{W}_{-1}\left(1+2\log(2/\alpha)+2\log\zeta(s)+2s(1-\log(2s))+2s\log(2s+\log n)\right)}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{2\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}}\right]$ (71)
with $n>n_{\min}$ where $n_{\min}=\operatorname*{polylog}(1/\alpha)$ is the largest $n$ such that the denominator above is nonpositive, form an $(1-\alpha)$-CS for $\mu$ over $\mathcal{N}$.
When $n$ is fixed, the growth rate of the CS as $\alpha\to 0$ is $\tilde{\mathcal{O}}\left(\sqrt{\exp(\frac{2(1+\eta)}{n-1}\log\frac{1}{\alpha})}\right)=\tilde{\mathcal{O}}(\alpha^{-\frac{1+\eta}{n-1}})$, depending on the stitching parameter $\eta>0$ chosen. This is less desirable compared to two of the $\mathcal{O}(\alpha^{-1/n})$-rate t-confidence sequences we derived in this paper; see Table 1. When $\alpha$ is fixed, the shrinkage rate of the CS as $n\to\infty$ is the optimal iterated logarithm rate $\sqrt{n^{-1}\log\log n}$; however, such asymptotic benefit, overshadowed by the nonasymptotic price of large constants, does not manifest even with very large values of $n$, as is shown in the simulation Figure 1. Furthermore, the plug-in approach does not give a test process: though both of the two steps are based on some supermartingales, no union bound argument conjoins them, unlike the case of confidence sequences. In practice, therefore, the plug-in approach by Jourdan et al. [2023] is less appealing than our universal inference and scale-invariant likelihood ratio mixture.
A.2 A Sequential Test for Median by Howard and Ramdas [2022]
Since Gaussian distributions are symmetric, any confidence sequence for the median is a confidence sequence for the mean as well. We quote the following recent result from Howard and Ramdas [2022, Theorem 1]:
Proposition A.4 (Confidence Sequence for Median).
Let
$f_{n}=0.75\sqrt{\ell_{n}}+0.8\ell_{n}\quad\ell_{n}=\frac{1.4\log\log 2.1n+\log(10/\alpha)}{n}.$ (72)
For any continuous distribution $P$ on $\mathbb{R}$,
$\operatorname{CI}_{n}^{\mathsf{md}}=[\widehat{Q}_{n}(0.5-f_{n}),\widehat{Q}_{n}^{-}(0.5+f_{n})]$ (73)
forms a $(1-\alpha)$ confidence sequence for the median of $P$; therefore also a $(1-\alpha)$ confidence sequence for $\mu$ over $\mathcal{N}$.
Here, the empirical quantiles $\widehat{Q}_{n}$ and $\widehat{Q}_{n}^{-}$ are defined as as
$\widehat{Q}_{n}(p)=\sup\{x:\widehat{F}_{n}(x)\leqslant p\},\quad\widehat{Q}_{n}^{-}(p)=\sup\{x:\widehat{F}_{n}(x)<p\}.$ (74)
When $n$ is fixed and $\alpha\to 0$, we note, the CS eventually becomes the entire $\mathbb{R}$ as long as $f_{n}>0.5$, which is true when $\log(10/\alpha)>n/10$. If, on the other hand, letting $\alpha_{\star}(n)$ be the value of $\alpha$ such that $f_{n}$ is exactly $0.5$, then, as $\alpha\downarrow\alpha_{\star}(n)=\exp(-\widetilde{\Theta}(n))$,
$|\operatorname{CI}_{n}^{\mathsf{md}}|\asymp 2\sigma\Phi^{-1}(0.5+f_{n})=\widetilde{\Theta}(\sqrt{\log(1/(0.5-f_{n}))})=\widetilde{\Theta}\left(\sqrt{\log(1/(\alpha-\alpha_{\star}(n)))}\right)$ (75)
under $N_{\mu,\sigma^{2}}$ and large $n$ (where $\widetilde{\Theta}$ hides $\log\log(1/(\alpha-\alpha_{\star}(n)))$-order terms).
When $\alpha$ is fixed, the shrinkage rate of the CS as $n\to\infty$ also enjoys the optimal iterated logarithm rate $\sqrt{n^{-1}\log\log n}$ due to the stitching technique it employs; to be precise,
$\lim_{n\to\infty}\frac{n\sigma^{-2}|\operatorname{CI}_{n}^{\mathsf{md}}|^{2}-4.5\pi\log(10/\alpha)}{\log\log n}=6.3\pi,\quad\text{almost surely under $N_{\mu,\sigma^{2}}$},$ (76)
indicating that when $n$ and $1/\alpha$ grows simultaneously, in the rate of $n=\widetilde{\Theta}(\log(1/\alpha))$ (where $\widetilde{\Theta}$ hides $\log\log(1/\alpha)$-order terms), the width of the CS stays bounded, a property satisfied by all of those $\mathcal{O}(\alpha^{-1/n})$-width CSs mentioned so far as well.
Again, as is shown in Figures 1 and 2, this asymptotically optimal rate does not translate into practical advantages for sample size $<10000$.
Proposition A.4 is based on the following test supermartingale, a Chernoff bound on i.i.d. Bernoulli indicator random variables due to Howard et al. [2020, Fact 1 (b)].
Proposition A.5 (One-Sided Test Supermartingale for Median).
For any constant $\lambda\in\mathbb{R}$, the process
$B_{n}^{\lambda}=\exp\left(\lambda\sum_{i=1}^{n}\mathbbmss{1}_{\{X_{i}>0\}}-\frac{n\lambda}{2}-n\log\left(\frac{\exp(\lambda/2)+\exp(-\lambda/2)}{2}\right)\right)$ (77)
is a supermartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ for the set of all continuous distributions with median 0; in particular, for $\mathcal{N}_{\mu=0}$.
It is straightforward to see that the growth of $\{B_{n}^{\lambda}\}$ under some $N_{\mu,\sigma^{2}}$ has the following growth index:
$\lim_{n\to\infty}\frac{\log(B_{n}^{\lambda})}{n}=\lambda(\Phi(\mu/\sigma)-1/2)-\log\left(\frac{\exp(\lambda/2)+\exp(-\lambda/2)}{2}\right)\quad\text{almost surely}.$ (78)
Here, $\Phi$ is the cumulative distribution function of $N_{0,1}$. As is shown in Figure 4, this is a very small growth index compared to the optimal $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, and is powerful only against the alternative $\mathcal{N}_{\mu>0}$ when $\lambda>0$. Therefore, Howard and Ramdas [2022] consider different choices of $\lambda$ for different epochs of time (“stitching”) to obtain Proposition A.4, which is still much looser than many of the other CSs we derive. Another way to gain universal power against all $\mathcal{N}_{\mu\neq 0}$ is, again, via the method of mixture, and Howard and Ramdas [2022] consider a beta-binomial mixture distribution on $\lambda$. Below, we let $\operatorname{B}(a,b)$ denote the beta function.
Proposition A.6 (Beta-Binomial Mixture Test Supermartingale for Median).
Let $T_{n}=\sum_{i=1}^{n}\mathbbmss{1}_{\{X_{i}>0\}}-n/2$. For any $a,b>0$, the process
$B_{n}^{(a,b)}=\frac{\operatorname{B}(a+n/2+T_{n},b+n/2-T_{n})}{(1/2)^{n}\operatorname{B}(a,b)}$ (79)
is a supermartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ for the set of all continuous distributions with median 0; in particular, for $\mathcal{N}_{\mu=0}$. Under $N_{\mu,\sigma^{2}}$, it has the following growth index:
$\lim_{n\to\infty}\frac{\log(B_{n}^{(a,b)})}{n}=\log\left(\Phi(\mu/\sigma)^{\Phi(\mu/\sigma)}\Phi(-\mu/\sigma)^{\Phi(-\mu/\sigma)}\right)-\log(1/2).$ (80)
Figure 4: The growth indices of the one-sided median test supermartingale (78) and mixture test supermartingale (80) under different choices of $\lambda$ and alternative effect sizes $\mu/\sigma$. The “upper bound” refers to the optimal growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
This statement, proved in Section C.5, shows that the mixture does gain power over all $\mu\neq 0$ as is depicted in Figure 4. However, compared to the optimal growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, it falls short in power especially when the effect size $\mu/\sigma$ is large.
Remark A.7.
The convergence (80) also holds in expectation (because the sequence is uniformly integrable). Following Vovk and Wang [2022, Section 5], we may define a concept of asymptotic relative efficiency (ARE) of the test process $\{B_{n}^{(a,b)}\}$ here, as the ratio between the second derivatives at 0 of the growth index curves of the upper bound and $\{B_{n}^{(a,b)}\}$ in Figure 4:
$\operatorname{ARE}(B_{n}^{(a,b)})=\lim_{\theta\to 0}\frac{\frac{1}{2}\log(1+\theta^{2})}{\log\left(\Phi(\theta)^{\Phi(\theta)}\Phi(-\theta)^{\Phi(-\theta)}\right)-\log(1/2)}$ (81)
$=$ $\frac{\frac{\mathrm{d}^{2}}{\mathrm{d}\theta^{2}}\frac{1}{2}\log(1+\theta^{2})|_{\theta=0}}{\frac{\mathrm{d}^{2}}{\mathrm{d}\theta^{2}}\log\left(\Phi(\theta)^{\Phi(\theta)}\Phi(-\theta)^{\Phi(-\theta)}\right)|_{\theta=0}}=\sqrt{\pi/8}\approx 0.627.$ (82)
This quantity bears the following meaning: fix any $\beta>0$ and define by $n(\mu,\sigma^{2})$ the sample size when $\mathrm{E}_{N_{\mu,\sigma^{2}}}(\log B_{n}^{(a,b)})$ first reaches $\beta$, $n^{*}(\mu,\sigma^{2})$ the sample size when the upper bound of expected log value of e-processes, $\frac{n}{2}\log(1+\mu^{2}/\sigma^{2})$, first reaches $\beta$ under $N_{\mu,\sigma^{2}}$. Then, $\operatorname{ARE}(B_{n}^{(a,b)})=\lim_{\mu/\sigma\to 0}\frac{n(\mu,\sigma^{2})}{n^{*}(\mu,\sigma^{2})}$. In comparison, we note that in the fixed-time setting the classical Pitman ARE of the sign test (which Proposition A.6 sequentializes) relative to the optimal t-test is $2/\pi\approx 0.637$ in the Gaussian case [Sprent and Smeeton, 2007, p.74]. Note that the fixed-time sign test is exact, whereas the sequential Proposition A.6 is conservative.
Appendix B Universal Inference Z-Tests
B.1 Two-Sided Z-Test
Recall from Corollary 3.1 that, for any $\mu$, $\sigma$ and their point estimators $\{\widetilde{\mu}_{n}\}$, $\{\widetilde{\sigma}_{n}\}$ adapted to the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$, the process
$\ell_{n}^{\mu,\sigma^{2}}=\frac{\prod_{i=1}^{n}p_{\widetilde{\mu}_{i-1},\widetilde{\sigma}^{2}_{i-1}}(X_{i})}{\prod_{i=1}^{n}p_{\mu,\sigma^{2}}(X_{i})}=\frac{\sigma^{n}}{\prod_{i=1}^{n}\widetilde{\sigma}_{i-1}}\exp\left\{\sum_{i=1}^{n}\left(\frac{(X_{i}-\mu)^{2}}{2\sigma^{2}}-\frac{(X_{i}-\widetilde{\mu}_{i-1})^{2}}{2\widetilde{\sigma}^{2}_{i-1}}\right)\right\}$ (83)
is a martingale for $N_{\mu,\sigma^{2}}$ on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
For the two-sided Z-test, we consider the point null $N_{0,\sigma^{2}}$ and the alternative $\mathcal{N}_{\mu\neq 0,\sigma^{2}}$. In this case, the variance $\sigma^{2}$ being a known quantity, the plug-in estimates $\{\widetilde{\sigma}_{n}^{2}\}$ can be taken simply as $\sigma^{2}$. We have the following exponential growth asymptotic which is slightly faster than the $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ rate we see so far for t-tests. Both rates are optimal in their respective regimes due to our calculation in Section 6.3.
Proposition B.1 (Asymptotic behavior of the universal inference Z-test e-process).
Under any $P=N_{\mu,\sigma^{2}}$, suppose there is a $\gamma>0$ such that $\{\widetilde{\mu}_{n}\}$ converges to $\mu$ in $L^{2}$ with rate $\mathrm{E}_{P}(|\widetilde{\mu}_{n}-\mu|^{2})\lesssim n^{-\gamma}$, and $\widetilde{\sigma}_{n}^{2}=\sigma^{2}$ for all $n$. Then,
$\lim_{n\to\infty}\frac{\log\ell_{n}^{0,\sigma^{2}}}{n}=\frac{\mu^{2}}{2\sigma^{2}}\quad\text{almost surely}.$ (84)
Consequently, $\{\ell_{n}^{0,\sigma^{2}}\}$ diverges almost surely to $\ell^{0,\sigma^{2}}_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu\neq 0}$.
Proof.
$\frac{1}{n}\log\ell_{n}^{0,\sigma^{2}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{X_{i}^{2}}{2\sigma^{2}}-\frac{(X_{i}-\widetilde{\mu}_{i-1})^{2}}{2\sigma^{2}}\right)=\frac{1}{2\sigma^{2}}\overline{X_{n}^{2}}-\frac{(X_{i}-\widetilde{\mu}_{i-1})^{2}}{2\sigma^{2}}.$ (85)
Repeating the steps in the Proof of Proposition 3.3 in Appendix C, we see that the first of these terms converges a.s. to $(\mu^{2}+\sigma^{2})/2\sigma^{2}$, and the last to $1/2$. ∎
B.2 One-Sided Z-Test
An infimum over $\mu\leqslant 0$ yields the following one-sided Z-test.
Proposition B.2 (Universal inference one-sided Z-test e-process).
For any $\mu$, $\sigma$ and their point estimators $\{\widetilde{\mu}_{n}\}$, $\{\widetilde{\sigma}_{n}\}$ adapted to the canonical filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$, the process
$\ell_{n}^{\sigma^{2}-}=\inf_{\mu\leqslant 0}\ell_{n}^{\mu,\sigma^{2}}=\frac{\sigma^{n}}{\prod_{i=1}^{n}\widetilde{\sigma}_{i-1}}\exp\left\{\sum_{i=1}^{n}\left(\frac{V_{n}-n(\widehat{\mu}_{n}\wedge 0)^{2}}{2\sigma^{2}}-\frac{(X_{i}-\widetilde{\mu}_{i-1})^{2}}{2\widetilde{\sigma}^{2}_{i-1}}\right)\right\}$ (86)
is an e-process for $N_{\mu\leqslant 0,\sigma^{2}}$ on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
The corresponding asymptotic can be similarly derived so we omit its proof.
Proposition B.3 (Asymptotic behavior of universal inference one-sided Z-test e-process).
Under any $P=N_{\mu,\sigma^{2}}$, suppose there is a $\gamma>0$ such that $\{\widetilde{\mu}_{n}\}$ converges to $\mu$ in $L^{2}$ with rate $\mathrm{E}_{P}(|\widetilde{\mu}_{n}-\mu|^{2})\lesssim n^{-\gamma}$, and $\widetilde{\sigma}_{n}^{2}=\sigma^{2}$ for all $n$. Then,
$\lim_{n\to\infty}\frac{\log\ell_{n}^{\sigma^{2}-}}{n}=\frac{(\mu\vee 0)^{2}}{2\sigma^{2}}\quad\text{almost surely}.$ (87)
Consequently, $\{\ell_{n}^{\sigma^{2}-}\}$ diverges almost surely to $\ell^{\sigma^{2}-}_{\infty}=\infty$ exponentially fast under $\mathcal{N}_{\mu>0}$.
Appendix C Omitted Proofs
C.1 Proofs for Likelihood Ratio Martingales
Proof of Lemma 2.4.
It suffices to prove that
$\mathrm{E}_{P}\left(\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(X_{n})\middle|X_{1},\dots,X_{n-1}\right)=1.$ (88)
For any $B\in\mathcal{B}(\mathbb{R})$, by the Doob–Dynkin factorization lemma (e.g. Klenke [2013, Corollary 1.97]), since $\omega\mapsto Q_{n}(\omega,B)$ is $\mathcal{F}_{n-1}$ measurable, there is a measurable $q(\cdot,B):\mathbb{R}^{n-1}\to\mathbb{R}$ such that $q(X_{1},\dots,X_{n-1},B)=Q_{n}(\omega,B)$.
Let $A\subseteq\mathbb{R}^{n-1}$ be measurable. Then,
$\mathrm{E}_{P}\left(\mathbbmss{1}_{\{(X_{1},\dots,X_{n-1})\in A\}}\cdot\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(X_{n})\right)$ (89)
$=$ $\int_{A\times\mathbb{R}}\frac{\mathrm{d}q_{n}(x_{1},\dots,x_{n-1},\cdot)}{\mathrm{d}P}(x_{n})\,\mathrm{d}P^{\otimes n}(x_{1},\dots,x_{n})$ (90)
$=$ $\int_{A}\left\{\int_{\mathbb{R}}\frac{\mathrm{d}q_{n}(x_{1},\dots,x_{n-1},\cdot)}{\mathrm{d}P}(x_{n})\,\mathrm{d}P(x_{n})\right\}\,\mathrm{d}P^{\otimes(n-1)}(x_{1},\dots,x_{n-1})$ (91)
$=$ $\int_{A}q_{n}(x_{1},\dots,x_{n-1},\mathbb{R})\,\mathrm{d}P^{\otimes(n-1)}(x_{1},\dots,x_{n-1})$ (92)
$=$ $\int_{A}\mathrm{d}P^{\otimes(n-1)}(x_{1},\dots,x_{n-1})=P^{\otimes(n-1)}(A).$ (93)
Since $A$ is arbitrary, we conclude that
$\mathrm{E}_{P}\left(\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(X_{n})\middle|X_{1},\dots,X_{n-1}\right)=1,$ (94)
which yields the desired result. ∎
Proof of Lemma 2.5.
Let $A\subseteq\mathbb{R}^{n-1}$ be measurable. Then,
[…]
And
[…]
Since $A$ is arbitrary, we conclude that
[…]
C.2 Proofs for Universal Inference t-Tests
Proof of Theorem 3.2.
For each $\sigma>0$ consider the likelihood ratio martingale under $\mathcal{N}(\mu,\sigma^{2})$
$M_{n}^{\mu,\sigma^{2}}=\frac{\prod_{i=1}^{n}p_{\widetilde{\mu}_{i-1},\widetilde{\sigma}^{2}_{i-1}}(X_{i})}{\prod_{i=1}^{n}p_{\mu,\sigma^{2}}(X_{i})}.$ (100)
The following process is an e-process under any $\mathcal{N}(\mu,\cdot)$
$R_{n}^{\mu}=\frac{\prod_{i=1}^{n}p_{\widetilde{\mu}_{i-1},\widetilde{\sigma}^{2}_{i-1}}(X_{i})}{\sup_{s>0}\prod_{i=1}^{n}p_{\mu,s^{2}}(X_{i})}=\inf_{s>0}M_{n}^{\mu,s^{2}}.$ (101)
Note that
$\sup_{\sigma>0}\operatorname*{\mathbb{P}}_{\mathcal{N}(\mu,\sigma^{2})}[\exists n,R_{n}^{\mu}\geqslant 1/\alpha]\leqslant\operatorname*{\mathbb{P}}_{\mathcal{N}(\mu,\sigma^{2})}[\exists n,M_{n}^{\mu,\sigma^{2}}\geqslant 1/\alpha]\leqslant\alpha.$ (102)
Since $s=\sqrt{\frac{\sum(X_{i}-\mu)^{2}}{n}}$ maximizes the denominator of $R_{n}^{\mu}$, it has a closed-form expression:
$R_{n}^{\mu}=\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\sup_{s>0}\prod\frac{1}{s}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\mu}{s}\right)^{2}\right\}}=\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\left(\frac{\sum(X_{i}-\mu)^{2}}{n}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}.$ (103)
Hence due to (102), with probability at least $1-\alpha$, for all $n$,
$\frac{\sum(X_{i}-\mu)^{2}}{n}\leqslant$$\left\{\ \frac{1}{\alpha}\exp\left\{-\frac{n}{2}\right\}\prod{\widetilde{\sigma}_{i-1}}\exp\left\{\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}\ \right\}^{2/n}$ (104)
$=$ $\frac{1}{\alpha^{2/n}\mathrm{e}}\prod_{i=1}^{n}{\widetilde{\sigma}_{i-1}^{2/n}}\exp\left\{\frac{1}{n}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}$ (105)
$=$ $\underbrace{\frac{1}{\alpha^{2/n}\mathrm{e}}\exp\left\{\frac{\sum_{i=1}^{n}\log\widetilde{\sigma}_{i-1}^{2}+\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}}{n}\right\}}_{T_{n}}$ (106)
Recall that the random variable on the RHS is $T_{n}$. The UI-CS then reads
$\operatorname{CI}_{n}=\left[\widehat{\mu}_{n}\pm\sqrt{\widehat{\mu}_{n}^{2}-\overline{X^{2}_{n}}+T_{n}}\right].$ (107)
Prood of Proposition 3.3.
Note that
$\frac{1}{n}\log R_{n}=\frac{1}{2}\log{\overline{X^{2}_{n}}}+\frac{1}{2}-\frac{\sum_{i=1}^{n}\log\widetilde{\sigma}_{i-1}}{n}-\frac{\sum_{i=1}^{n}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}}{2n}.$ (108)
The first three terms converge a.s. to $\frac{1}{2}\log(\mu^{2}+\sigma^{2})$, $1/2$, and $\log\sigma$. It remains to show that the last term above converges a.s. to $1/2$.
Define $A_{n}=\sum_{i=1}^{n}\frac{1}{i}\left|\left(\frac{X_{i}-\mu}{\widetilde{\sigma}_{i-1}}\right)^{2}-\left(\frac{X_{i}-\mu}{\sigma}\right)^{2}\right|$. Then,
$\mathrm{E}_{P}(A_{n})=\sum_{i=1}^{n}\frac{1}{i}\mathrm{E}_{P}\left(|X_{i}-\mu|^{2}\cdot|\widetilde{\sigma}_{i-1}^{-2}-\sigma^{-2}|\right)$ (109)
$\leqslant$ $\sum_{i=1}^{n}\frac{1}{i}\mathrm{E}_{P}\left(\frac{(X_{i}-\mu)^{4}i^{-\gamma/2}+(\widetilde{\sigma}_{i-1}^{-2}-\sigma^{-2})^{2}i^{\gamma/2}}{2}\right)$ (110)
$=$ $\frac{1}{2}\sum_{i=1}^{n}\frac{1}{i}\left(\mathrm{E}_{P}(X_{i}-\mu)^{4}\cdot i^{-\gamma/2}+\mathrm{E}_{P}(\widetilde{\sigma}_{i-1}^{-2}-\sigma^{-2})^{2}\cdot i^{\gamma/2}\right)$ (111)
$=$ $\frac{1}{2}\sum_{i=1}^{n}\frac{1}{i}(\mathcal{O}(1)\cdot i^{-\gamma/2}+\mathcal{O}(i^{-\gamma})\cdot i^{\gamma/2})=\frac{1}{2}\sum_{i=1}^{n}\mathcal{O}(i^{-1-\gamma/2})=\mathcal{O}(1).$ (112)
So there exists a constant $K>0$ such that $\mathrm{E}_{P}(A_{n})\leqslant K$ for all $n$. By Markov’s inequality and Fatou’s lemma,
$P(\lim_{n\to\infty}A_{n}\geqslant a)\leqslant\frac{\mathrm{E}_{P}(\lim_{n\to\infty}A_{n})}{a}\leqslant\frac{\liminf_{n\to\infty}\mathrm{E}_{P}(A_{n})}{a}\leqslant\frac{K}{a},$ (113)
which implies $P(\lim_{n\to\infty}A_{n}<\infty)=1$. By Kronecker’s lemma, this implies that
$\lim_{n\to\infty}\frac{\sum_{i=1}^{n}\left\{\left(\frac{X_{i}-\mu}{\widetilde{\sigma}_{i-1}}\right)^{2}-\left(\frac{X_{i}-\mu}{\sigma}\right)^{2}\right\}}{n}=0\quad\text{almost surely.}$. (114)
We now define $B_{n}=\sum_{i=1}^{n}\frac{1}{i}\left|\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}-\left(\frac{X_{i}-\mu}{\widetilde{\sigma}_{i-1}}\right)^{2}\right|$. Without loss of generality assume $\sup_{n\geqslant 0}\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-6})<\infty$ (otherwise, if only $\sup_{n\geqslant n_{0}}\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-6})<\infty$ holds, use summation $\sum_{i=n_{0}+1}^{n}$ instead). Then,
$\mathrm{E}_{P}(B_{n})=\sum_{i=1}^{n}\frac{1}{i}\mathrm{E}_{P}\left(\widetilde{\sigma}^{-2}_{i-1}\cdot|2X_{i}-\widetilde{\mu}_{i-1}-\mu|\cdot|\widetilde{\mu}_{i-1}-\mu_{1}|\right)$ (115)
$\leqslant\sum_{i=1}^{n}\frac{1}{i}\mathrm{E}_{P}\left(\frac{\widetilde{\sigma}^{-6}_{i-1}i^{-\gamma/4}+|2X_{i}-\widetilde{\mu}_{i-1}-\mu|^{3}i^{-\gamma/4}+|\widetilde{\mu}_{i-1}-\mu|^{3}i^{\gamma/2}}{3}\right)$ (116)
$=\frac{1}{3}\sum_{i=1}^{n}\frac{1}{i}\left(\mathrm{E}_{P}(\widetilde{\sigma}^{-6}_{i-1})\cdot i^{-\gamma/4}+\mathrm{E}_{P}(|2X_{i}-\widetilde{\mu}_{i-1}-\mu|^{3})\cdot i^{-\gamma/4}+\mathrm{E}_{P}(|\widetilde{\mu}_{i-1}-\mu|^{3})\cdot i^{\gamma/2}\right)$ (117)
$=\frac{1}{3}\sum_{i=1}^{n}\frac{1}{i}\left(\mathcal{O}(1)\cdot i^{-\gamma/4}+\mathcal{O}(1)\cdot i^{-\gamma/4}+\mathcal{O}(i^{-\gamma})\cdot i^{\gamma/2}\right)=\frac{1}{3}\sum_{i=1}^{n}\mathcal{O}(i^{-1-\gamma/4})=\mathcal{O}(1).$ (118)
Similarly by Markov’s inequality, Fatou’s and Kronecker’s lemma,
$\lim_{n\to\infty}\frac{\sum_{i=1}^{n}\left\{\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}-\left(\frac{X_{i}-\mu}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{n}=0\quad\text{almost surely.}$. (119)
Finally, strong law of large numbers implies that
$\lim_{n\to\infty}\frac{\sum_{i=1}^{n}\left(\frac{X_{i}-\mu}{\sigma}\right)^{2}}{n}=1\quad\text{almost surely.}$. (120)
Combining (114), (119), and (120), we see that
$\lim_{n\to\infty}\frac{\sum_{i=1}^{n}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}}{n}=1\quad\text{almost surely.}$. (121)
This concludes the proof. ∎
Proof of Theorem 3.4.
The following process is an e-process under any $\mathcal{N}(\mu,\sigma^{2})$, ($\mu\leqslant 0$ and $\sigma>0$).
$\frac{\prod_{i=1}^{n}p_{\widetilde{\mu}_{i-1},\widetilde{\sigma}^{2}_{i-1}}(X_{i})}{\sup_{\begin{subarray}{c}m\leqslant 0\\ s>0\end{subarray}}\prod_{i=1}^{n}p_{m,s^{2}}(X_{i})}=\inf_{\begin{subarray}{c}m\leqslant 0\\ s>0\end{subarray}}M_{n}^{m,s^{2}}.$ (122)
We have
$\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\sup_{m\leqslant 0}\sup_{s>0}\prod\frac{1}{s}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-m}{s}\right)^{2}\right\}}=\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\sup_{m\leqslant 0}\left(\frac{\sum(X_{i}-m)^{2}}{n}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}$ (123)
$=\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\sup_{m\leqslant 0}\left(m^{2}-2\widehat{\mu}_{n}m+\overline{X^{2}_{n}}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}$ (124)
The denominator is maximized by $m=\widehat{\mu}_{n}\wedge 0$. Therefore
$\dots$=⋯ = $\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\left((\widehat{\mu}_{n}\wedge 0-\widehat{\mu}_{n})^{2}+\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}=\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\left((0\vee\widehat{\mu}_{n})^{2}+\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}$ (125)
$=$ $\frac{\prod\frac{1}{\widetilde{\sigma}_{i-1}}\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\left(\overline{X^{2}_{n}}-(\widehat{\mu}_{n}\wedge 0)^{2}\right)^{-n/2}\exp\left\{-\frac{n}{2}\right\}}=R_{n}^{-}.$ (126)
Proof of Proposition 3.5.
The four terms of
$\frac{1}{n}\log R_{n}^{-}=\frac{1}{2}\log\left(\overline{X^{2}_{n}}-(\widehat{\mu}_{n}\wedge 0)^{2}\right)+\frac{1}{2}-\frac{\sum_{i=1}^{n}\log\widetilde{\sigma}_{i-1}}{n}-\frac{\sum_{i=1}^{n}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}}{2n}$ (127)
converges a.s. to $\frac{1}{2}\log(\mu^{2}+\sigma^{2}-(\mu\wedge 0)^{2})$, $1/2$, $\log\sigma$, and $1/2$ respectively, concluding the proof. ∎
C.3 Proofs for Scale Invariant t-Tests
Proof of Lemma 4.4.
Let $Q^{+}_{n}$ be the distribution of $(X_{1},X_{2},\dots,X_{n})$ conditioned on $\{X_{1}>0\}$. Define the constant $p_{>0}=\mathbb{P}[X_{1}>0]=P_{\mu,\sigma}\{(0,\infty)\}$. Then,
$\frac{\mathrm{d}Q^{+}_{n}}{\mathrm{d}\eta^{\otimes n}}(x_{1},x_{2},\dots,x_{n})=p_{>0}^{-1}\cdot\mathbbmss{1}(x_{1}>0)\cdot\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}-\mu))\right\},$ (128)
meaning that for any measurable $A\subseteq\mathbb{R^{+}}\times\mathbb{R}^{n-1}$,
$Q_{n}^{+}\{A\}=\int_{A}p_{>0}^{-1}\cdot\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}-\mu))\right\}\eta(\mathrm{d}x_{1})\eta(\mathrm{d}x_{2})\dots\eta(\mathrm{d}x_{n}).$ (129)
Now recall $V_{n}:(x_{1},x_{2},\dots,x_{n})\mapsto(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|)$. Let $Q_{n}^{*+}$ be the push-forward measure of $Q_{n}^{+}$ by $V_{n}$, a measure on $\{1\}\times\mathbb{R}^{n-1}$ (which is the conditional distribution of $(X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|)$ given $\{X_{1}>0\}$). Suppose the set $V_{n}(A)=\{1\}\times B$. We have the following change of variables (letting $(x_{1}^{*},\dots,x_{n}^{*})=V_{n}(x_{1},\dots,x_{n})$, so $x_{1}^{*}=1$ below),
$Q_{n}^{*+}\{V_{n}(A)\}=Q_{n}^{+}\{A\}=\int_{A}p_{>0}^{-1}\cdot\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}-\mu))\right\}\eta(\mathrm{d}x_{1})\eta(\mathrm{d}x_{2})\dots\eta(\mathrm{d}x_{n})$ (130)
$=p_{>0}^{-1}\cdot\int_{x_{1}>0}\eta(\mathrm{d}x_{1})\int_{B}\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}^{*}|x_{1}|-\mu))\right\}\eta(\mathrm{d}x_{2}^{*})\dots\eta(\mathrm{d}x_{n}^{*})\left|\frac{\mathrm{d}(x_{2},\dots,x_{n})}{\mathrm{d}(x_{2}^{*},\dots,x_{n}^{*})}\right|$ (131)
$=p_{>0}^{-1}\cdot\int_{x_{1}>0}\eta(\mathrm{d}x_{1})\int_{B}\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}^{*}|x_{1}|-\mu))\right\}\eta(\mathrm{d}x_{2}^{*})\dots\eta(\mathrm{d}x_{n}^{*})\begin{vmatrix}\frac{1}{|x_{1}|}&0&\dots&0\\ 0&\frac{1}{|x_{1}|}&&0\\ \vdots&&\ddots&0\\ 0&\dots&0&\frac{1}{|x_{1}|}\end{vmatrix}^{-1}$ (132)
$=p_{>0}^{-1}\cdot\int_{x_{1}>0}\eta(\mathrm{d}x_{1})\int_{B}\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}^{*}x_{1}-\mu))\right\}x_{1}^{n-1}\eta(\mathrm{d}x_{2}^{*})\dots\eta(\mathrm{d}x_{n}^{*}).$ (133)
We thus conclude that (letting $\delta_{1}$ be the Dirac point mass on 1),
$\frac{\mathrm{d}Q_{n}^{*+}}{\delta_{1}\otimes\eta^{\otimes(n-1)}}(x_{1}^{*},x_{2}^{*},\dots,x_{n}^{*})=p_{>0}^{-1}\cdot\int_{x_{1}>0}\prod_{i=1}^{n}\left\{\sigma^{-1}g(\sigma^{-1}(x_{i}^{*}x_{1}-\mu))\right\}x_{1}^{n-1}\eta(\mathrm{d}x_{1}).$ (134)
We can simplify the integral above by substituting $\tau=\sigma/x_{1}$, obtaining,
$\frac{\mathrm{d}Q_{n}^{*+}}{\delta_{1}\otimes\eta^{\otimes(n-1)}}(x_{1}^{*},x_{2}^{*},\dots,x_{n}^{*})$ (135)
$=$ $p_{>0}^{-1}\cdot\int_{\tau>0}\left\{\sigma^{-1}g(1/\tau-\sigma^{-1}\mu)\right\}\left\{\prod_{i=2}^{n}\sigma^{-1}g(x_{i}^{*}/\tau-\sigma^{-1}\mu))\right\}(\sigma/\tau)^{n-1}\sigma(1/\tau^{2})\eta(\mathrm{d}\tau)$ (136)
$=$ $p_{>0}^{-1}\cdot\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}^{*}/\tau-\mu/\sigma))\right\}(1/\tau)\mathrm{d}\tau.$ (137)
Similarly, let $Q_{n}^{*-}$ be the conditional distribution of $(X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|)$ given $\{X_{1}<0\}$). We have
$\frac{\mathrm{d}Q_{n}^{*-}}{\delta_{-1}\otimes\eta^{\otimes(n-1)}}(x_{1}^{*},x_{2}^{*},\dots,x_{n}^{*})=(1-p_{>0})^{-1}\cdot\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}^{*}/\tau-\mu/\sigma))\right\}(1/\tau)\mathrm{d}\tau.$ (138)
Finally, the measure $Q_{n}^{*}=\mathbbmss{1}_{\{x_{1}>0\}}\cdot p_{>0}\cdot Q_{n}^{*+}+\mathbbmss{1}_{\{x_{1}<0\}}\cdot(1-p_{>0})\cdot Q_{n}^{*-}$. So we see that
$\frac{\mathrm{d}Q_{n}}{\mathrm{d}\eta^{\langle n-1\rangle}}(x_{1}^{*},\dots,x_{n}^{*})=\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}^{*}/\tau-\mu/\sigma))\right\}\frac{\mathrm{d}\tau}{\tau}$ (139)
holds for both $x_{1}^{*}=\pm 1$. This concludes the proof. ∎
From now, let us denote
$p^{*}_{\theta}(x_{1}^{*},\dots,x_{n}^{*})=\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}^{*}/\tau-\theta))\right\}(1/\tau)\mathrm{d}\tau,$ (140)
i.e. the density $\frac{\mathrm{d}Q_{n}}{\mathrm{d}\eta^{\langle n-1\rangle}}$ when $P_{\mu,\sigma^{2}}$ satisfies $\mu/\sigma=\theta$. Here is a property of $p_{\theta}^{*}$.
Proposition C.1.
Let $\lambda>0$ be any constant. Then, $p^{*}_{\theta}(\lambda x_{1}^{*},\dots,\lambda x_{n}^{*})=\lambda^{-n}p^{*}_{\theta}(x_{1}^{*},\dots,x_{n}^{*})$. In particular, taking $\lambda=|x_{1}|$, we have $p^{*}_{\theta}(x_{1},\dots,x_{n})=|x_{1}|^{-n}p^{*}_{\theta}(x_{1}^{*},\dots,x_{n}^{*})$.
Proof of Proposition C.1.
$p^{*}_{\theta}(\lambda x_{1}^{*},\dots,\lambda x_{n}^{*})$ (141)
$=$ $\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(\lambda x_{i}^{*}/\tau-\theta))\right\}(1/\tau)\mathrm{d}\tau$ (142)
$=$ $\int_{\gamma=\lambda^{-1}\tau>0}\left\{\prod_{i=1}^{n}(1/\lambda\gamma)g(x_{i}^{*}/\gamma-\theta))\right\}(1/\lambda\gamma)\lambda\mathrm{d}\gamma$ (143)
$=$ $\lambda^{-n}p^{*}_{\theta}(x_{1}^{*},\dots,x_{n}^{*}).$ (144)
Proof of Lemma 4.5.
Denote $X_{i}/|X_{1}|$ by $X_{i}^{*}$.
[…]
for any $\mu/\sigma=\theta$ and $\mu_{0}/\sigma_{0}=\theta_{0}$. Now it follows from Lemma 2.5 that $\{h_{n}(\theta;\theta_{0})\}$ is an NM for $P_{\mu_{0},\sigma^{2}_{0}}$ on the canonical filtration generated by $\{X_{n}^{*}\}$, which according to Proposition 4.3 is just $\{\mathcal{F}_{n}^{*}\}$. ∎
Proof of Corollary 4.6.
Using $g(x)=\frac{1}{\sqrt{2\pi}}\exp(-x^{2}/2)$ in (140),
$p^{*}_{\theta}(x_{1},\dots,x_{n})=\int_{\tau>0}\left\{\prod_{i=1}^{n}(1/\tau)g(x_{i}/\tau-\theta))\right\}(1/\tau)\mathrm{d}\tau$ (146)
$=$ $\frac{1}{(2\pi)^{n/2}}\int_{\tau>0}\left\{\prod_{i=1}^{n}\mathrm{e}^{-(x_{i}/\tau-\theta)^{2}/2}\right\}\frac{1}{\tau^{n+1}}\mathrm{d}\tau$ (147)
$=$ $\frac{\mathrm{e}^{-n\theta^{2}/2}}{(2\pi)^{n/2}}\int_{\tau>0}\exp\left\{-\frac{\sum_{i=1}^{n}{x_{i}}^{2}}{2\tau^{2}}+\frac{\theta\sum_{i=1}^{n}x_{i}}{\tau}\right\}\frac{1}{\tau^{n+1}}\mathrm{d}\tau.$ (148)
Therefore, applying Lemma 4.5 with $\theta_{0}=0$, we have
$h_{n}(\theta;0)=\mathrm{e}^{-n\theta^{2}/2}\frac{\int_{\tau>0}\exp\left\{-\frac{\sum_{i=1}^{n}{X_{i}}^{2}}{2\tau^{2}}+\frac{\theta\sum_{i=1}^{n}X_{i}}{\tau}\right\}\frac{1}{\tau^{n+1}}\mathrm{d}\tau}{\int_{\tau>0}\exp\left\{-\frac{\sum_{i=1}^{n}{X_{i}}^{2}}{2\tau^{2}}\right\}\frac{1}{\tau^{n+1}}\mathrm{d}\tau}$ (149)
(the denominator being a generalized Gaussian integral) (150)
$=$ $\frac{\mathrm{e}^{-n\theta^{2}/2}}{\Gamma(n/2)}\int_{y>0}y^{n/2-1}\exp\left\{-y+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}y,$ (151)
coinciding with $h_{\theta,n}$. ∎
Proof of Theorem 4.7 (as well as Theorem 4.1).
Putting a flat, improper prior $\mathrm{d}\pi(\theta)=\mathrm{d}\theta$ over $\mathbb{R}$, we have the following improperly mixed, extended nonnegative supermartingale,
$H_{n}=\int h_{\theta,n}\pi(\mathrm{d}\theta)$ (152)
$=$ $\int_{\theta}\frac{\mathrm{e}^{-n\theta^{2}/2}}{\Gamma(n/2)}\left(\int_{y>0}y^{n/2-1}\exp\left\{-y+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}y\right)\cdot\mathrm{d}\theta$ (153)
$=$ $\frac{1}{\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\int_{\theta}\exp\left\{-\frac{n}{2}\theta^{2}+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}\theta\right)\mathrm{d}y$ (154)
$=$ $\frac{1}{\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\sqrt{\frac{2\pi}{n}}\exp\left\{\frac{\left(\sum_{i=1}^{n}X_{i}\right)^{2}\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}{2n}\right\}\right)\mathrm{d}y$ (155)
$=$ $\frac{1}{\Gamma(n/2)}\sqrt{\frac{2\pi}{n}}\int_{y>0}y^{n/2-1}\exp\left\{-\left(1-\frac{S_{n}^{2}}{nV_{n}}\right)y\right\}\mathrm{d}y$ (156)
$=$ $\frac{1}{\Gamma(n/2)}\sqrt{\frac{2\pi}{n}}\Gamma(n/2)\left(1-\frac{S_{n}^{2}}{nV_{n}}\right)^{-n/2}=\sqrt{\frac{2\pi}{n}}\left(1-\frac{S_{n}^{2}}{nV_{n}}\right)^{-n/2}$ (157)
$=$ $\sqrt{\frac{2\pi}{n}}\left(\frac{nV_{n}}{nV_{n}-S_{n}^{2}}\right)^{n/2}$ (158)
Note that when $n=1$, $nV_{n}-S_{n}^{2}=0$, and $H_{0}$ is understood to be +$\infty$+ ∞. Note that
$t_{n-1}\sim\frac{\frac{1}{n}S_{n}}{\frac{1}{\sqrt{n}}\sqrt{\frac{1}{n-1}(V_{n}-\frac{1}{n}S_{n}^{2})}}=\sqrt{n-1}\sqrt{\frac{S_{n}^{2}}{nV_{n}-S_{n}^{2}}}=:T_{n-1}$ (159)
Then $H_{n}=\sqrt{\frac{2\pi}{n}}\left(\frac{T_{n-1}^{2}}{n-1}+1\right)^{n/2}$.
In terms of CS, every $X_{i}$ should be seen as $X_{i}-0=X_{i}-\mu$, and hence
$H_{n}^{\mu}=\sqrt{\frac{2\pi}{n}}\left(\frac{n\sum(X_{i}-\mu)^{2}}{n\sum(X_{i}-\mu)^{2}-(\sum(X_{i}-\mu))^{2}}\right)^{n/2}$ (160)
is also an extended nonnegative supermartingale.
The extended Ville’s inequality, applied on the process $\sqrt{m/2\pi}H_{n}^{\mu}$, reads,
With probability at most
$P=\mathbb{P}\left[\frac{T_{m-1}^{2}}{m-1}+1\geqslant\eta\right]+\eta^{-m/2}\mathbb{E}\left[\mathbbmss{1}_{\left\{\frac{T_{m-1}^{2}}{m-1}+1<\eta\right\}}\left(\frac{T_{m-1}^{2}}{m-1}+1\right)^{m/2}\right],$ (161)
there exists $n\geqslant m$,
$\sqrt{\frac{m}{n}}\left(\frac{n\sum(X_{i}-\mu)^{2}}{n\sum(X_{i}-\mu)^{2}-(\sum(X_{i}-\mu))^{2}}\right)^{n/2}\geqslant\eta^{m/2}.$ (162)
Define $a=\sqrt{(m-1)(\eta-1)}$. Then $\eta^{m/2}=(1+a^{2}/(m-1))^{m/2}=h(a)$ and
$P$ $=\mathbb{P}[|T_{m-1}|\geqslant a]+h(a)^{-1}\mathbb{E}[\mathbbmss{1}_{\{|T_{m-1}|<a\}}h(T_{m-1})]$ (163)
$=2\mathbb{P}[T_{m-1}>a]+2\frac{\mathbb{E}[\mathbbmss{1}_{\{0\leqslant T_{m-1}<a\}}h(T_{m-1})]}{h(a)}$ (164)
Let $f_{m-1}$, $F_{m-1}$ be the density and CDF of $T_{m-1}$. The expectation $\mathbb{E}[\mathbbmss{1}_{\{0\leqslant T_{m-1}<a\}}h(T_{m-1})]$ must be smaller than $\frac{a}{a-a^{\prime}}\mathbb{E}[\mathbbmss{1}_{\{a^{\prime}\leqslant T_{m-1}<a\}}h(T_{m-1})]$, for all $a^{\prime}\in(0,a)$. Note that
$\lim_{a^{\prime}\to a}\frac{a}{a-a^{\prime}}\mathbb{E}[\mathbbmss{1}_{\{a^{\prime}\leqslant T_{m-1}<a\}}h(T_{m-1})]=ah(a)f_{m-1}(a).$ (165)
Hence
$P\leqslant 2(1-F_{m-1}(a))+2af_{m-1}(a).$ (166)
This concludes the missing proof of Lai’s t-CS. ∎
Proof of Theorem 4.9.
Let us put a Gaussian prior on $\theta$,
$\frac{\mathrm{d}\pi(\theta)}{\mathrm{d}\theta}=\frac{c}{\sqrt{2\pi}}\mathrm{e}^{-c^{2}\theta^{2}/2},$ (167)
and define the martingale
$G_{n}^{(c)}=\int h_{\theta,n}\pi(\mathrm{d}\theta)$ (168)
$=$ $\int_{\theta}\frac{\mathrm{e}^{-n\theta^{2}/2}}{\Gamma(n/2)}\left(\int_{y>0}y^{n/2-1}\exp\left\{-y+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}y\right)\cdot\pi(\mathrm{d}\theta)$ (169)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\int_{\theta}\exp\left\{-\frac{n+c^{2}}{2}\theta^{2}+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}\theta\right)\mathrm{d}y$ (170)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\exp\left\{\frac{\left(\sum_{i=1}^{n}X_{i}\right)^{2}\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}{2(n+c^{2})}\right\}\right)\mathrm{d}y$ (171)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\int_{y>0}y^{n/2-1}\exp\left\{-\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)y\right\}\mathrm{d}y$ (172)
$=$ $\frac{c}{\Gamma(n/2)}\sqrt{\frac{1}{n+c^{2}}}\Gamma(n/2)\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}.$ (173)
This is a valid test martingale.
While this is a test martingale for $\mathcal{N}_{\mu=0}$, a shifting argument ($X_{i}\to X_{i}-\mu_{0}$) yields that
$G_{n}^{(c,\mu_{0})}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1-\frac{(S_{n}-n\mu_{0})^{2}}{(n+c^{2})(V_{n}-2S_{n}\mu_{0}+n\mu_{0}^{2})}\right)^{-n/2}$ (174)
is a test martingale for $\mathcal{N}_{\mu=\mu_{0}}$. Applying Ville’s inequality, with probability at least $1-\alpha$, for all $n$,
$\frac{nc^{2}\mu_{0}^{2}-2c^{2}S_{n}\mu_{0}+(n+c^{2})V_{n}-S_{n}^{2}}{(n+c^{2})(n\mu_{0}^{2}-2S_{n}\mu_{0}+V_{n})}$ $\geqslant\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n},$ (175)
$\frac{nc^{2}(\mu_{0}-\widehat{\mu}_{n})^{2}+(n+c^{2})(V_{n}-S_{n}^{2}/n)}{(n+c^{2})(n(\mu_{0}-\widehat{\mu}_{n})^{2}+V_{n}-S_{n}^{2}/n)}$ $\geqslant\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n},$ (176)
$\left\{\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})n-nc^{2}\right\}(\mu_{0}-\widehat{\mu}_{n})^{2}$$\leqslant\left\{(n+c^{2})\left(1-\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}\right)\right\}(V_{n}-S_{n}^{2}/n),$ (177)
$(\mu_{0}-\widehat{\mu}_{n})^{2}$ $\leqslant\frac{(n+c^{2})\left(1-\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}\right)}{\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}}\left(\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)$ (178)
This concludes the proof. ∎
Proof of Proposition 4.8.
First, under $N_{\mu,\sigma^{2}}$, note that $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu}{\sqrt{nV_{n}-S_{n}^{2}}}$ and $v_{n}=V_{n}/n-S_{n}^{2}/n^{2}$. We replace $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu_{1}}{\sqrt{v_{n}}}$ in (34) to have
$H_{n}=\sqrt{\frac{2\pi}{n}}\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}\right)^{n/2},$ (179)
$\frac{\log H_{n}}{n}=\frac{\log(2\pi/n)}{2n}+\frac{1}{2}\log\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}\right).$ (180)
It is well known $v_{n}$ converges to $\sigma^{2}$ almost surely. Therefore $\frac{T_{n-1}}{\sqrt{n-1}}=\frac{\widehat{\mu}_{n}-\mu_{1}}{\sqrt{v_{n}}}$ converges almost surely to 0. These imply that $\frac{\log H_{n}}{n}$ converges almost surely to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
Second, under any $P\in\mathcal{N}_{\mu=0}$, let $\delta>0$. By (34) we have
$P(H_{n}>\delta)=P\Bigg{(}T_{n-1}^{2}>\underbrace{\left(\left(\frac{\delta^{2}n}{2\pi}\right)^{1/n}-1\right)(n-1)}_{\gtrsim\log n}\Bigg{)}\lesssim P(T_{1}^{2}>\log n).$ (181)
The underbraced term grows logarithmically because $n^{1/n}-1=\frac{\log n}{n}+\mathcal{O}(n^{-2})$. The $\lesssim$” holds because the tail of Student’s t-distribution strictly decreases when the DOF increases. So $P(H_{n}>\delta)$ goes to 0 as $n\to\infty$, concluding that $\{H_{n}\}$ converges to 0 in probability. By Wang and Ramdas [2023a, Proposition A.14], an ENSM converges almost surely, so $\{H_{n}\}$ converges to 0 almost surely. The proof is complete. ∎
Proof of Proposition 4.10.
First, under $N_{\mu,\sigma^{2}}$, we replace in (42) $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}$,
$G_{n}^{(c)}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}}+{c^{2}}}\right)^{n/2},$ (182)
$\frac{\log G_{n}^{(c)}}{n}=\frac{\log(\frac{c^{2}}{n+c^{2}})}{2n}+\frac{1}{2}\log\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}}+{c^{2}}}\right).$ (183)
As in the proof of Proposition 4.8, $\frac{T_{n-1}}{\sqrt{n-1}}$ converges to 0 and $v_{n}$ to $\sigma^{2}$ almost surely. Therefore $\frac{\log G_{n}^{(c)}}{n}$ converges a.s. to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
Second, under any $P\in\mathcal{N}_{\mu=0}$, let $\delta>0$. By (42) we have
$P(G_{n}^{(c)}>\delta)=P\Bigg{(}\frac{n}{\frac{(n+c^{2})(n-1)}{T_{n-1}^{2}}+{c^{2}}}>\underbrace{\left(\frac{(n+c^{2})\delta^{2}}{c^{2}}\right)^{1/n}-1}_{\gtrsim\frac{\log n}{n}}\Bigg{)}\lesssim P(T_{1}^{2}>\log n),$ (184)
which converges to 0. So $\{G_{n}^{(c)}\}$ converges to 0 in probability. It thus also converges to 0 almost surely due to Doob’s martingale convergence theorem, concluding the proof. ∎
Proof of Theorem 4.11.
Let us put a half Gaussian prior on $\theta$,
$\frac{\mathrm{d}\pi^{+}(\theta)}{\mathrm{d}\theta}=\mathbbmss{1}_{\theta\geqslant 0}\frac{c}{\sqrt{2\pi}}\mathrm{e}^{-c^{2}\theta^{2}/2},$ (185)
and define the martingale
$\int h_{\theta,n}\pi^{+}(\mathrm{d}\theta)$ (186)
$=$ $\int_{\theta>0}\frac{\mathrm{e}^{-n\theta^{2}/2}}{\Gamma(n/2)}\left(\int_{y>0}y^{n/2-1}\exp\left\{-y+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}y\right)\cdot\pi(\mathrm{d}\theta)$ (187)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\int_{\theta>0}\exp\left\{-\frac{n+c^{2}}{2}\theta^{2}+\theta\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}\right\}\mathrm{d}\theta\right)\mathrm{d}y$ (188)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\int_{y>0}y^{n/2-1}\mathrm{e}^{-y}\left(\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\exp\left\{\frac{\left(\sum_{i=1}^{n}X_{i}\right)^{2}\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}{2(n+c^{2})}\right\}\Phi\left(\frac{\left(\sum_{i=1}^{n}X_{i}\right)\sqrt{\frac{2y}{\sum_{i=1}^{n}{X_{i}}^{2}}}}{\sqrt{n+c^{2}}}\right)\right)\mathrm{d}y$ (189)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\int_{y>0}y^{n/2-1}\exp\left\{-\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)y\right\}\Phi\left(\frac{S_{n}\sqrt{2y}}{\sqrt{(n+c^{2})V_{n}}}\right)\mathrm{d}y$ (190)
[…]
$\geqslant$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\int_{y>0}y^{n/2-1}\exp\left\{-\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)y\right\}\left(1-\exp\left(-\frac{(S_{n}\vee 0)^{2}y}{(n+c^{2})V_{n}}\right)\right)\mathrm{d}y$ (192)
$=$ $\frac{c}{\sqrt{2\pi}\Gamma(n/2)}\sqrt{\frac{\pi}{\frac{n+c^{2}}{2}}}\bigg{\{}\int_{y>0}y^{n/2-1}\exp\left\{-\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)y\right\}\mathrm{d}y$ (193)
$-\int_{y>0}y^{n/2-1}\exp\left(\frac{(S_{n}^{2}-(S_{n}\vee 0)^{2})y}{(n+c^{2})V_{n}}-1\right)\mathrm{d}y\bigg{\}}$ (194)
$=$ $\frac{c}{\Gamma(n/2)}\sqrt{\frac{1}{n+c^{2}}}\left\{\Gamma(n/2)\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}-\Gamma(n/2)\left(1-\frac{(S_{n}\wedge 0)^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}\right\}$ (195)
$=$ $\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}-\left(1-\frac{(S_{n}\wedge 0)^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}\right)=:G^{(c-)}_{n}/2.$ (196)
Since $\pi^{+}$ is of total measure $1/2$, $G^{c+}_{n}$ is an e-process for $\mathcal{N}_{\mu=0}$. ∎
C.4 Proofs for Optimality
Proof of Theorem 6.1.
Fix an $L\in\mathcal{L}$, $\sigma>0$, and $\alpha>0$. Let
$\mu=\sigma\sqrt{(1-(1-3\alpha)^{2})^{-2/n}-1}.$ (197)
First, consider the distribution $N_{0,\mu^{2}+\sigma^{2}}$. We have
$N_{0,\mu^{2}+\sigma^{2}}(0<L(X_{1},\dots,X_{n}))\leqslant\alpha.$ (198)
Next, consider the distribution $N_{\mu,\sigma^{2}}$. Since $D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}^{\otimes n}\|N_{0,\mu^{2}+\sigma^{2}}^{\otimes n})=\frac{n}{2}\log(1+\mu^{2}/\sigma^{2})$, it follows from Bretagnolle-Huber inequality that
$\operatorname{TV}(N_{\mu,\sigma^{2}}^{\otimes n},N_{0,\mu^{2}+\sigma^{2}}^{\otimes n})\leqslant\sqrt{1-\exp(-D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}^{\otimes n}\|N_{0,\mu^{2}+\sigma^{2}}^{\otimes n}))}=\sqrt{1-(1+\mu^{2}/\sigma^{2})^{-n/2}}.$ (199)
Therefore,
$N_{\mu,\sigma^{2}}(0<L(X_{1},\dots,X_{n}))\leqslant\alpha+\sqrt{1-(1+\mu^{2}/\sigma^{2})^{-n/2}}=\alpha+1-3\alpha=1-2\alpha.$ (200)
By the definition of $W_{2\alpha}(L,P)$,
$N_{\mu,\sigma^{2}}(\mu-L(X_{1},\dots,X_{n})<W_{2\alpha}(L,P))=1-2\alpha,$ (201)
indicating that $W_{2\alpha}(L,N_{\mu,\sigma^{2}})\geqslant\mu=\sigma\sqrt{(1-(1-3\alpha)^{2})^{-2/n}-1}$. Since this holds for any lower confidence bound $L\in\mathcal{L}_{\alpha,n}$, we see that
$M^{-}_{\alpha,n}=\inf_{L\in\mathcal{L}_{\alpha,n}}\sup_{P\in\mathcal{N}}\sigma^{-1}(P)\cdot W_{2\alpha}(L,P)\geqslant$ $\sqrt{(1-(1-3\alpha)^{2})^{-2/n}-1}$ (202)
$=$ $\sqrt{(6\alpha-9\alpha^{2})^{-2/n}-1}.$ (203)
Proof of Corollary 6.2.
Fix a $\operatorname{CI}=[L(X_{1},\dots,X_{n}),U(X_{1},\dots,X_{n})]\in\mathcal{I}_{\alpha,n}$ and consider any $P\in\mathcal{N}$. First, since $L\in\mathcal{L}_{\alpha,n}$, we have
$P(\mu(P)-L(X_{1},\dots,X_{n})\geqslant W_{2\alpha}(L,P))=2\alpha.$ (204)
Next, since $[L(X_{1},\dots,X_{n}),U(X_{1},\dots,X_{n})]$ is a $(1-\alpha)$-CI,
$P(\mu(P)\leqslant U(X_{1},\dots,X_{n}))\geqslant 1-\alpha.$ (205)
By a union bound,
$P(U(X_{1},\dots,X_{n})-L(X_{1},\dots,X_{n})\geqslant W_{2\alpha}(L,P))\geqslant\alpha,$ (206)
implying that $W_{\alpha}(\operatorname{CI},P)\geqslant W_{2\alpha}(L,P)$. Therefore $M_{\alpha,n}\geqslant M_{\alpha,n}^{-}$ by taking a minimax. ∎
Proof of Proposition 6.4.
Let $P=N_{0,\sigma^{2}_{0}}$. Then
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|P)=\frac{1}{2}(\log\sigma_{0}^{2}-\log\sigma^{2})+\frac{\sigma^{2}+\mu^{2}}{2\sigma_{0}^{2}}-\frac{1}{2}.$ (207)
As a function of $\sigma_{0}^{2}$, the above takes minimum when $\sigma_{0}^{2}=\sigma^{2}+\mu^{2}$, in which case the minimum is $D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}+\mu^{2}})=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
Under any $P_{0}=N_{0,\sigma_{0}^{2}}\in\mathcal{P}$:
$\mathrm{E}_{P_{0}}\left(\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})\right)=\int\frac{\sqrt{\sigma^{2}+\mu^{2}}}{\sigma\sigma_{0}\sqrt{2\pi}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}+\frac{x^{2}}{2(\sigma^{2}+\mu^{2})}-\frac{x^{2}}{2\sigma_{0}^{2}}\right)\mathrm{d}x$ (208)
$=$ $\frac{\sqrt{\sigma^{2}+\mu^{2}}}{\sigma\sigma_{0}\sqrt{2\pi}}\int\exp\left(-\frac{\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2})}{2\sigma^{2}\sigma_{0}^{2}(\sigma^{2}+\mu^{2})}x^{2}+\frac{\mu}{\sigma^{2}}x-\frac{\mu^{2}}{2\sigma^{2}}\right)\mathrm{d}x$ (209)
$=$ $\frac{\sqrt{\sigma^{2}+\mu^{2}}}{\sigma\sigma_{0}\sqrt{2\pi}}\cdot\sqrt{\frac{\pi}{\frac{\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2})}{2\sigma^{2}\sigma_{0}^{2}(\sigma^{2}+\mu^{2})}}}\exp\left(\frac{\mu^{2}/\sigma^{4}}{4\frac{\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2})}{2\sigma^{2}\sigma_{0}^{2}(\sigma^{2}+\mu^{2})}}-\frac{\mu^{2}}{2\sigma^{2}}\right)$ (210)
$=$ $\frac{\sigma^{2}+\mu^{2}}{\sqrt{\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2})}}\exp\left(\frac{\mu^{2}(\sigma_{0}^{2}-(\sigma^{2}+\mu^{2}))}{2(\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2}))}\right)$ (211)
letting $z=\sqrt{\mu^{2}\sigma_{0}^{2}+\sigma^{2}(\sigma^{2}+\mu^{2})}$ and $y=\sigma^{2}+\mu^{2}$ (212)
$=$ $\frac{y}{z}\exp\left(\frac{z^{2}-y^{2}}{2z^{2}}\right)=\sqrt{\frac{y^{2}}{z^{2}}\exp\left(1-\frac{y^{2}}{z^{2}}\right)}\leqslant 1.$ (213)
This concludes that the likelihood ratio is a supermartingale for $\mathcal{N}_{\mu=0}$. The rest of the statement follows from the basic facts we mentioned about numeraires and the reference therein. ∎
C.5 Proofs for Additional t-Tests
Proof of Proposition A.6.
Following Howard et al. [2021, Proposition 7], let $p=\frac{\exp(\lambda/2)}{\exp(\lambda/2)+\exp(-\lambda/2)}$. We can now write the process $B_{n}^{\lambda}$ as
$B_{n}^{\lambda}=\frac{p^{n/2+T_{n}}(1-p)^{n/2-T_{n}}}{(1/2)^{n}}.$ (214)
Now, integrating the supermartingale above with $p\sim\operatorname{Beta}(a,b)$,
$B_{n}^{(a,b)}=\int_{0}^{1}\frac{p^{n/2+T_{n}}(1-p)^{n/2-T_{n}}}{(1/2)^{n/2+T_{n}}(1/2)^{n/2-T_{n}}}\cdot\frac{p^{a-1}(1-p)^{b-1}}{\operatorname{B}(a,b)}\mathrm{d}p=\frac{\operatorname{B}(a+n/2+T_{n},b+n/2-T_{n})}{(1/2)^{n}\operatorname{B}(a,b)}$ (215)
forms a supermartingale as well.
Next, due to the facts
  1. 1.
    $\lim_{n\to\infty}\frac{\log\operatorname{B}(\alpha n,\beta n)}{n}=\log\frac{\alpha^{\alpha}\beta^{\beta}}{(\alpha+\beta)^{\alpha+\beta}}$ for any $\alpha,\beta>0$; and
  2. 2.
    $T_{n}/n\to\Phi(\mu/\sigma)-1/2$ almost surely,
we can see that the test supermartingale $\{B_{n}^{(a,b)}\}$ under some $N_{\mu,\sigma^{2}}$ has the following growth index:
$\lim_{n\to\infty}\frac{\log(B_{n}^{(a,b)})}{n}=\log\left(\Phi(\mu/\sigma)^{\Phi(\mu/\sigma)}\Phi(-\mu/\sigma)^{\Phi(-\mu/\sigma)}\right)-\log(1/2).$ (216)